### v1

# NeuronSeek: On Stability and Expressivity of Task-driven Neurons

Abstract
Drawing inspiration from our human brain that designs different neurons for different tasks, recent advances in deep learning have explored modifying a network’s neurons to develop so-called task-driven neurons. Prototyping task-driven neurons (referred to as NeuronSeek) employs symbolic regression (SR) to discover the optimal neuron formulation and construct a network from these optimized neurons. Along this direction, this work replaces symbolic regression with tensor decomposition (TD) to discover optimal neuronal formulations, offering enhanced stability and faster convergence. Furthermore, we establish theoretical guarantees that modifying the aggregation functions with common activation functions can empower a network with a fixed number of parameters to approximate any continuous function with an arbitrarily small error, providing a rigorous mathematical foundation for the NeuronSeek framework. Extensive empirical evaluations demonstrate that our NeuronSeek-TD framework not only achieves superior stability, but also is competitive relative to the state-of-the-art models across diverse benchmarks. The code is available at https://github.com/HanyuPei22/NeuronSeek.
Index Terms:
Neuronal diversity, tensor decomposition, task-driven neuron, super-expressive
I Introduction
Deep learning has achieved significant success in a wide range of fields [1, 2]. Credits for these accomplishments largely go to the so-called scaling law, wherein fundamental computational units such as convolutions and attention mechanisms are designed and scaled per an optimized architecture. Prominent examples include ResNet [3], Transformer [4], and Mamba [5]. Recently, motivated by neuroscience’s paramount contributions to AI, an emerging field called NeuroAI has garnered significant attention, which draws upon the principles of biological circuits in the human brain to catalyze the next revolution in AI [6]. The ambitious goal of NeuroAI is rooted in the belief that the human brain, as one of the most intelligent systems, inherently possesses the capacity to address complex challenges in AI development [7]. Hence, it can always serve as a valuable source of inspiration for guiding practitioners, despite the correspondence between biological and artificial neural networks sometimes being implicit.
Following NeuroAI, let us analyze the scaling law through the lens of brain computation. It is seen that the human brain generates complex intellectual behaviors through the cooperative activity of billions of mutually connected neurons with varied morphologies and functionalities [8]. This suggests that the brain simultaneously benefits from scale at the macroscopic level and neuronal diversity at the microscopic level. The latter is a natural consequence of stem cells’ directed programming in order to facilitate the task-specific information processing. Inspired by this observation, extensive research has shown that incorporating neuronal diversity into artificial neural networks can markedly enhance their capabilities [7]. For instance, researchers have transitioned beyond traditional inner-product neurons by integrating diverse nonlinear aggregation functions, such as quadratic neurons [9], polynomial neurons [10], and dendritic neurons [11]. Such networks have achieved superior performance in tasks like image recognition and generation, underscoring the promise and practicality of innovating a neural network at the neuronal level.
Along this direction, driven by the task-specific nature of neurons in our brain, recent research [12] introduced a systematic framework for task-driven neurons. Hereafter, we refer to prototyping task-driven neurons as NeuronSeek. This approach assumes that there is no single type of neurons that can perform universally well. Therefore, it enables neurons to be tailored for specific tasks through a two-stage process. As shown in Figure 1, in the first stage, the vectorized symbolic regression (SR)—an extension of symbolic regression [13, 14] that enforces uniform operations for all variables—identifies the optimal homogeneous neuronal formula from data. Unlike traditional linear regression and symbolic regression, the vectorized SR searches for coefficients and homogeneous formulas at the same time, which can facilitate parallel computing. In the second stage, the derived formula serves as the neuron’s aggregation function, with learnable coefficients. Notably, activation functions remain unmodified in task-driven neurons. At the neuronal level, task-driven neurons are supposed to have superior performance than those one-size-fits-all units do, as it imbues the prior information pertinent to the task. More favorably, task-driven neurons can still be connected into a network to fully leverage the power of connectionism. Hereafter, we refer to task-driven neurons using symbolic regression as NeuronSeek-SR. NeuronSeek considers in tandem neuronal importance and scale. From design to implementation, it systematically integrates three cornerstone AI paradigms: behaviorism (the genetic algorithm runs symbolic regression), symbolism (symbolic regression distills a formula from data), and connectionism.
Figure 1: Task-driven neuron based on the vectorized symbolic regression (NeuronSeek-SR).
Despite this initial success, the following two major challenges remain unresolved in prototyping task-driven neurons.
  • Non-deterministic and unstable convergence. While symbolic regression employs modified genetic programming (GP) to discover polynomial expressions [15, 16], its tree-based evolutionary approach struggles to converge in high-dimensional space. The resulting formulas demonstrate acute sensitivity to certain hyperparameters, such as initialization conditions, population size, and mutation rate. This may confuse users about whether the identified formula is truly suitable for the task or due to the random perturbation, therefore hurting the legibility of task-driven neurons.
  • Necessity of task-driven neurons. In the realm of deep learning theory, it was shown [17, 18, 19] that certain activation functions can empower a neural network using a fixed number of neurons to approximate any continuous function with an arbitrarily small error. These functions are termed “super-expressive” activation functions. Such a unique and desirable property allows a network to achieve a precise approximation without increasing structural complexity. Given the tremendous theoretical advantages of adjusting activation functions, it is necessary to address the following issue: Can the super-expressive property be achieved via revising the aggregation function while retaining common activation functions?
In this study, we address the above two issues satisfactorily. On one hand, in response to the instability of NeuronSeek-SR, we propose a tensor decomposition (TD) method referred to as NeuronSeek-TD. This approach begins by assuming that the optimal representation of data is encapsulated by a high-order polynomial coupled with trigonometric functions. We construct the basic formulation and apply TD to optimize its low-rank structure and coefficients. Therefore, we reformulate the unstable formula search problem (NeuronSeek-SR) into a stable low-rank coefficient optimization task. To enhance robustness, we introduce the sparsity regularization in the decomposition process, automatically eliminating insignificant terms while ensuring the framework to identify the optimal formula. Additionally, rank regularization is employed to derive the simplest possible representation. The improved stability of our method stems from two key factors: i) Unlike conventional symbolic regression which suffers from complex hyperparameter tuning in genetic programming [20], NeuronSeek-TD renders significantly fewer tunable parameters. ii) It has been proven that tensor decomposition with rank regularization tends to yield a unique solution [21], thereby effectively addressing the inconsistency issues inherent in symbolic regression methods.
On the other hand, we close the theoretical deficit by showing that task-driven neurons that use common activation such as ReLU can also achieve the super-expressive property. Earlier theories like [17, 18, 19] first turn the approximation problem into the point-fitting problem, and then uses the dense trajectory to realize the super-expressive property. The key message is the existence of a one-dimensional dense trajectory to cover the space of interest. In this study, we highlight that dense trajectory can also be achieved by adjusting aggregation functions. Specifically, we integrate task-driven neurons into a discrete dynamical system defined by the layer-wise transformation $\mathbf{x}_{N}=T(\mathbf{x}_{N-1})$, where $T$ is the mapping performed by one layer of the task-driven network. There exists some initial point $\mathbf{x}_{0}$ that can be taken by $T$ to the neighborhood of any target point $\mathbf{x}$. To approximate $\mathbf{x}$, our construction does not need to increase the values of parameters in $T$ and the network, and we just compose the transformation module $T$ different times. This means that we realize “super-super-expressiveness”: not only the number of parameters but also the magnitudes remain fixed. In summary, our contributions are threefold:
  • We introduce tensor decomposition to prototype stable task-driven neurons by enhancing the stability of the process of finding a suitable formula for the given task. Our study is a major modification to NeuronSeek.
  • We theoretically show that task-driven neurons with common activation functions also enjoy the “super-super-expressive” property, which puts the task-driven neuron on a solid theoretical foundation. The novelty of our construction lies in that it fixes both the number and values of parameters, while the previous construction only fixes the number of parameters.
  • Systematic experiments on public datasets and real-world applications demonstrate that NeuronSeek-TD achieves not only superior performance but also stability compared to NeuronSeek-SR. Moreover, its performance is also competitive compared to other state-of-the-art baselines.
II Related Works
II-A Neuronal Diversity and Task-driven Neurons
Many groundbreaking AI advancements were inspired by the computing of biological neural systems. For instance, the neocognitron [22], a precursor to convolutional neural networks, is exactly a miniature of cortical visual circuits. The human brain is renowned for its extensive array of neurons, which vary significantly in morphology and functionality. This neuronal diversity underpins the brain’s intelligent behavior  [23, 24]. However, traditional artificial neural networks often prioritize the scaling law that designs a fundamental unit and replicates it across large architectures. This approach pays little attention to neuronal diversity. Recently, the idea of introducing neuronal diversity has naturally emerged, which mainly modifies the aggregation function instead of the activation function in constructing new neurons. This operation is well-grounded: Once an activation function is monotonic, a neuron’s decision boundary is exclusively shaped by its aggregation function. In particular, it replaces the inner product of a neuron with a nonlinear function such as a polynomial. Table I summarizes the recently proposed non-linear neurons. Notably, the complexity of neurons in [25, 26, 27] is of $\mathcal{O}(n^{2})$, which is much larger than the conventional neuron, while neurons in [28, 29, 9, 30] favorably enjoy the linear parametric complexity.
TABLE I: A summary of the recently-proposed neurons. $\sigma(\cdot)$ is the nonlinear activation function. $\odot$ denotes Hadamard product. $\mathbf{W}\in\mathbb{R}^{n\times n}$, $\mathbf{w_{i}}\in\mathbb{R}^{n\times 1}$, and the bias terms in these neurons are omitted for simplicity.
Works Formulations
Zoumponuris et al. (2017)  [25] $\mathbf{y}=\sigma(\mathbf{x}^{\top}\mathbf{W}\mathbf{x}+\mathbf{w}^{\top}\mathbf{x})$
Fan et al. (2018) [30] $\mathbf{y}=\sigma((\mathbf{w}_{1}^{\top}\mathbf{x})(\mathbf{w}_{2}^{\top}\mathbf{x})+\mathbf{w}_{3}^{\top}(\mathbf{x}\odot\mathbf{x}))$
Jiang et al. (2019) [26] $\mathbf{y}=\sigma(\mathbf{x}^{\top}\mathbf{W}\mathbf{x})$
Mantini & Shah (2021) [27] $\mathbf{y}=\sigma(\mathbf{w}^{\top}(\mathbf{x}\odot\mathbf{x}))$
Goyal et al. (2020) [28] $\mathbf{y}=\sigma(\mathbf{w}^{\top}(\mathbf{x}\odot\mathbf{x}))$
Bu & Karpante (2021) [29] $\mathbf{y}=\sigma((\mathbf{w}_{1}^{\top}\mathbf{x})(\mathbf{w}_{2}^{\top}\mathbf{x}))$
Xu et al. (2022) [9] $\mathbf{y}=\sigma((\mathbf{w}_{1}^{\top}\mathbf{x})(\mathbf{w}_{2}^{\top}\mathbf{x})+\mathbf{w}_{3}^{\top}\mathbf{x})$
Fan et al. (2024) [12] Task-driven polynomial
Following the way paved for task-specific neuronal design [12], this work constructs neurons by tensor decomposition. It simultaneously resolves theoretical deficits and improves the empirical stability of the original framework, representing a substantial advance in handling high-dimensional learning tasks like image recognition. While Chrysos et al. [31, 32, 33] focus on designing polynomial network architectures, our work differs by concentrating on task-specific neuronal construction and more diverse functional forms such as trigonometric functions ($\sin(\mathbf{x})$, $\cos(\mathbf{x})$).
II-B Symbolic Regression and Deep Learning
Symbolic regression (SR) is an algorithm that searches for formulas from data [13]. Recent research in SR includes learning partial differential equations (PDEs) from data [34, 35], improving the validity of SR on high-dimensional data [36], and invertible SR [37]. Besides, the Kolmogorov-Arnold Network (KAN) is an exemplary work that combines symbolic regression and deep learning [38]. KAN is inspired by the celebrated Kolmogorov-Arnold representation theorem, which states that any continuous multivariate function can be represented as a superposition of several continuous univariate functions. Such a decomposability lays KAN a solid theoretical foundation for function approximation.
Specifically, Kolmogorov’s theorem asserts that for any vector $\mathbf{x}=\left[x_{1},x_{2},\cdots,x_{d}\right]^{T}\in[0,1]^{d}$, a continuous function $f(\mathbf{x})$ can be expressed as
$f(x_{1},x_{2},\cdots,x_{d})=\sum_{i=0}^{2d}g_{i}\left(\sum_{j=1}^{d}h_{i,j}\left(x_{j}\right)\right),$ (1)
where $g_{i}$ and $h_{i,j}$ are continuous univariate functions. Building upon this theorem, KAN is a neural network to exploit this decomposition by learning $g_{i}$ and $h_{i,j}$, which can be regarded as a special symbolic regression. Currently, this KAN-type line has been successfully used in a myraid of domains, such as medical image analysis [39] and scientific discovery [40].
Moreover, deep learning is also applied to strengthen the performance of SR algorithms. Kamienny et al. [41] combined the Transformer and symbolic regression to enhance its searching ability. Kim et al. [42] integrated the neural network with symbolic regression, and proposed a framework similar to the equation learner network. SymbolicGPT, a Transformer-based language model for symbolic regression [43], is also shown to be a competent model compared with the existing models in terms of accuracy and run time.
To summarize, symbolism is one of the earliest and most influential paradigms in the history of AI [44, 45]. It is rooted in the idea that intelligence can be modeled through the manipulation of symbols—abstract representations of objects, concepts, or relationships in the real world. This paradigm assumes that reasoning and problem-solving can be achieved through symbolic manipulation. Symbolism and connectionism are highly complementary [46, 47]. Incorporating techniques in symbolism can effectively solve the intrinsic problems in connectionism such as efficiency, robustness, and forgetting. Our work synergizes symbolic regression and deep learning at the neuronal level, which is a novel practice in fusing symbolism and connectionism.
II-C Super-expressive Activation Functions
These days, super-expressive activation functions gain lots of traction in deep learning approximation theory [48]. Unlike the traditional activation functions, super-expressive activation functions can empower a network to achieve an arbitrarily small error with a fixed network width and depth. The idea of super-expressive activation function dates back to [49], while the activation function therein was not in closed form. Recently, several studies proposed explicit super-expressive activation functions [17, 18, 19], such as $\sin(x),\arcsin(x)$, and $x-\mathtt{floor}(x)$ with $x>0$. The key to achieving super-expressiveness lies in leveraging the dense orbit to solve the point-fitting problem $(n,f_{n})$, where $f_{n}\in[0,1],n=1,2,\cdots,N$. The fact that there exists $\theta$ such that $f_{n}=\theta/(\pi+n)-\mathtt{floor}(\theta/(\pi+n))$ was used in [18].
Our work shows that modifying the aggregation function can even endow a network with the super-super-expressive property. Particularly, our theoretical construction is much more efficient in memory. Previously, though the number of parameters like $\theta$ is fixed in [18], $\theta$ grows large when the approximation error goes low, and thus the essential network memory is still larger. In contrast, our construction maintains both a fixed number and magnitude of parameters, which brings a significant memory saving. Our work provides a strong theoretical foundation for task-driven neurons. Our proof is inspired by [50] that also composes a fixed block to do approximation. [50] uses the bit extract technique, while ours uses the chaotic theory.
Figure 2: The overall framework of the proposed method. In the first stage, the input data, regardless of tables and images, are flattened into a vector representation and processed using an initial formula for neuronal search. A stable formula is then generated through CP decomposition. In the second stage, the neuronal formula is parameterized and integrated into various neural network backbones for task-specific applications.
III Task-driven Neuron Based on Tensor Decomposition (NeuronSeek-TD)
III-A NeuronSeek-SR
In this section, we briefly illustrate how NeuronSeek-SR creates novel neurons. It follows a two-stage process, as illustrated in Figure 1. First, the vectorized SR identifies the optimal homogeneous neuronal formula by fitting task-specific data. Second, the derived formula is implemented as the neuron’s aggregation function with learnable coefficients.
III-A1 Vectorized SR
Conventional SR encodes a mathematical expression as a tree-structured directed graph and utilizes GP to navigate the solution space [15]. The GP optimization process employs biologically-inspired evolutionary operators, including: i) crossover operations that exchange subtrees between candidate solutions to maintain population diversity, and ii) mutation operations that introduce random modifications to explore novel regions of the function space. While this approach offers notable flexibility for symbolic optimization, it faces significant challenges such as computational complexity, formula heterogeneity, and overfitting when applied to high-dimensional problems. All these together are not ideal for the task of prototyping new neuronal formulations.
Unlike the conventional SR, which discovers heterogeneous formulas for each variable, the vectorized SR regularizes all input variables to learn the same formula [13, 14]. This single change brings huge advantages: i) Lower complexity. Homogeneous formulas generated by the vectorized SR have significantly fewer parameters, scaling linearly with the input dimensionality. ii) Faster computation and minor overfitting risks. By limiting the search space, the vectorized SR reduces the risk of overfitting and also makes the regression process faster and more computationally efficient, especially for high-dimensional data.
III-A2 Parameterization
The learned formulas from the vectorized SR are parameterized, i.e., making the coefficients trainable. This allows neurons to fine-tune their behavior in the context of a network. While task-based neurons alone may have limited approximation power, connecting them into a neural network framework will further amplify the power of neurons, thereby enabling efficient optimization and improving performance in handling complex tasks.
Despite the success of the vectorized SR, its built-in tree-based evolutionary algorithm still suffers from poor convergence in high-dimensional space. Consequently, the derived formulas exhibit significant sensitivity to key hyperparameters—including initialization conditions, population size, and mutation rate—often producing inconsistent solutions across different runs. This inherent instability in the NeuronSeek-SR framework stems from the intrinsic issues of genetic algorithms.
III-B NeuronSeek-TD
In this work, we propose NeuronSeek-TD, a faster and more stable alternative to conventional vectorized SR for neuronal formula discovery. We reasonably assert that the maximum required polynomial order for any given task is bounded by a constant $N$. By applying tensor decomposition to optimize coefficients, we reformulate the unstable formula search problem into a stable low-rank coefficient optimization task. Furthermore, we augment the polynomial structure through elementary functions, providing greater design flexibility and improved performance in high-dimensional pattern recognition tasks. The overall framework is illustrated in Figure 2, which comprises two stages: search for the neuronal formula and construct the task-driven neural network.
III-B1 Neuronal formula discovery
In the first stage, we adopt a polynomial model to identify the expressions that best fit the task-specific data. For consistent processing across modalities, both tabular and image inputs are flattened into a vector representation $\mathbf{z}\in\mathbb{R}^{d}$. Then, given the input data $\mathbf{z}$, the initial neuronal formula is defined as follows:
$f(\mathbf{z})$ $=C_{1}\cdot\beta_{j}+C_{2}\cdot\mathbf{z}^{\top}\mathbf{W}^{[1]}+C_{3}\cdot\mathbf{z}^{\top}\mathbf{W}^{[2]}\mathbf{z}$ (2)
$+C_{4}\cdot\mathbf{W}^{[3]}\times_{1}\mathbf{z}\times_{2}\mathbf{z}\times_{3}\mathbf{z}+\cdots+C_{N}\cdot\mathbf{W}^{[N]}\prod_{n=1}^{N}\times_{n}\mathbf{z}$
$+D\cdot\sin(\mathbf{z}),$
where $\mathbf{W}^{[n]}$ are weights and $C_{j}$ are coefficients. To put it in a more compact way, we have
$f(\mathbf{z})$ $=\sum_{n=1}^{N}\left(C_{n}\cdot\mathcal{W}^{[n]}\prod_{m=1}^{n}\mathbf{z}\right)+D\cdot\sin(\mathbf{z}),$ (3)
where the order $N$ and the rank are two hyperparameters to be set appropriately. The main idea of this method is to outer-product a series of rank-1 tensors to represent high-order tensors so that the computation cost is reduced significantly. The tensor decomposition process could be formulated as follows:
$\mathbf{W}^{[n]}=\sum_{r=1}^{R}\mathbf{a}_{r}^{(1)}\circ\mathbf{a}_{r}^{(2)}\circ\dots\circ\mathbf{a}_{r}^{(n)},$ (4)
where $R$ is decomposition rank, and $n$ is the order of the decomposed tensor $\mathbf{W}$. Here, we use COMP/PARAFAC (CP) decomposition to execute this process. CP decomposition has been widely adopted for modeling multilinear relationships in machine learning applications [51, 31]. In contrast to Tucker decomposition, CP decomposition provides both computational and structural advantages for our neuronal formula discovery framework. Computationally, CP decomposition requires only $O(nRd)$ parameters for an $n$-th order tensor of size $d^{n}$, compared to Tucker’s $O(R^{n}+nRd)$ complexity, enabling efficient handling of high-order polynomial terms. An example of CP decomposition is illustrated in Figure 4. While other tensor decomposition approaches could be explored (e.g., tensor ring decomposition, hierarchical Tucker decomposition) [52, 53], the efficiency and structural benefits of CP decomposition have already aligned well with the requirements of our polynomial modeling framework.
Figure 3: A rank-R CP decomposition of a third-order tensor.
Moreover, the incorporation of the trigonometric term $D\sin(\mathbf{z})$ serves a pivotal purpose. It extends the model’s capacity to capture periodic patterns and high-frequency components that intrinsically appear in many real-world datasets [54]. This is particularly crucial as pure polynomial expansions exhibit limited spectral sensitivity.
We then impose sparsity constraints through $L_{1}$ regularization on both the coefficients $C_{i}$ and weight tensors $\mathbf{W}^{[i]}$ to promote computational efficiency. Specifically, since each tensor $\mathbf{W}^{[i]}$ is decomposed using CP decomposition as $\mathbf{W}^{[i]}=\sum_{r=1}^{R}\mathbf{a}_{r}^{(1)}\circ\mathbf{a}_{r}^{(2)}\circ\dots\circ\mathbf{a}_{r}^{(i)}$, we apply sparsity regularization directly to the factor matrices. Thus, the complete optimization objective becomes
$\underset{\{\mathbf{a}_{r}^{(j)}\},C_{i}}{\min}\sum_{k=1}^{K}\mathcal{L}(f(\mathbf{x}_{k}),y_{k})+\lambda_{1}\|\mathbf{C}\|_{1}+\lambda_{2}\sum_{i=1}^{N}\sum_{r=1}^{R}\sum_{j=1}^{i}\|\mathbf{a}_{r}^{(j)}\|_{1},$ (5)
where $\mathcal{L}(\cdot)$ is the task-specific loss function, set to mean squared error (MSE) for regression and cross-entropy (CE) loss for classification. This structured sparsity approach encourages the discovery of interpretable and simplified polynomial formulas while maintaining computational efficiency through the low-rank tensor representation.
To further enhance the compactness of the discovered neuronal formula beyond the $L_{1}$ regularization’s sparsity induction, we introduce a post-training thresholding mechanism for explicit term selection. While the $L_{1}$ regularization effectively drives many coefficients toward zero, empirical observations reveal that a non-negligible portion of coefficients $C_{i}$ retain small but non-zero values, which cumulatively increase model complexity without substantial predictive contribution. The thresholding mechanism is a statistically grounded pruning strategy: After model convergence, we compute the mean $\mu_{|C|}$ and the standard deviation $\sigma_{|C|}$ of the absolute values of all coefficients $\{|C_{i}|\}_{i=1}^{N}$, then construct the final neuronal formula by retaining only terms whose coefficients satisfy $|C_{i}|\geq\tau$, where $\tau=\mu_{|C|}-\sigma_{|C|}$. This can further increase the stability of identified formulas.
Remark 1. Three key features of the proposed method warrant emphasis. i) The higher-order polynomial architecture systematically enables the learning of the polynomials. It is noticed that Eq. (2) is rather inclusive, e.g., it identifies multi-order variable interactions. This structural flexibility facilitates sophisticated nonlinear approximations, aligning the model more closely with the underlying data distribution. ii) The integration of trainable coefficients $\mathbf{C}$ with the $L_{1}$ regularization imposes selective sparsity constraints on the polynomial expansion. Specifically, the regularization term $\mathbf{C}$ induces sparsity by driving coefficients of less predictive terms toward zero, thereby automatically pruning nonessential higher-order components and yielding a parsimonious architecture. iii) The complementary $L_{1}$ regularization on the coefficients $\mathbf{W}$ enforces structured sparsity in the parameter space. This dual regularization strategy enhances feature selection while simultaneously improving computational efficiency—a critical advantage in high-dimensional contexts. Collectively, these design choices balance expressive capacity with compactness, ensuring robust performance without overparameterization.
III-B2 Task-driven network construction
In the second stage, the discovered neuronal formula is first parameterized into neuron-like structures. The selected polynomial and trigonometric terms are assigned trainable weights and integrated into neural network architectures. For tabular data processing, we employ fully connected neural networks. A polynomial with the maximum polynomial order of 3 is expressed as
${y}=\sigma(\mathbf{W}_{1}^{\top}\mathbf{z}+\mathbf{W}_{2}^{\top}\mathbf{z}\odot\mathbf{z}+\mathbf{W}_{3}^{\top}\mathbf{z}\odot\mathbf{z}\odot\mathbf{z}+{b}),$ (6)
with $\sigma$ being the ReLU activation function. If a specific order term is absent in the neuronal formula, its corresponding weight and bias parameters are initialized to zero.
Moreover, we extend our framework to convolutional neurons for images. As mentioned, the trigonometric term is incorporated to model periodic patterns and high-frequency components inherent in high-dimensional images. Consequently, the complete neuronal formula is structured as follows:
${y}=\sigma(\mathbf{W}_{1}^{\top}*\mathbf{z}+\mathbf{W}_{2}^{\top}*\mathbf{z}\odot\mathbf{z}+\mathbf{W}_{3}^{\top}*\mathbf{z}\odot\mathbf{z}\odot\mathbf{z}+\mathbf{W}_{4}*\sin(\mathbf{z})+{b}),$ (7)
where * denotes the convolution operation.
Finally, we introduce a simple initialization strategy to enhance training stability. Followed by [33], the weight matrix $\mathbf{W}_{1}$ is initialized as $\mathbf{W}_{1}\sim\mathcal{N}(0,\sigma^{2})$, where $\sigma=\sqrt{1/32(k_{n})}$, while all other weights are initialized to a tiny number such as 0.001. This strategy ensures that the network initially behaves like linear neurons and gradually learns higher-order terms during training. This strategy has been demonstrated the effectiveness in training quadratic neurons [55].
IV Super-expressive Property Theory of Task-driven Neuron
In this section, we formally prove that task-driven neurons that modify the aggregation function can also achieve the super-expressive property. Our theory is established based on the chaotic theory. Mathematically, we have the following theorem:
Theorem 1.
Let $f\in C([0,1])$ be a continuous function. Then, for any $\varepsilon>0$, there exists a function $h$ generated by a task-driven network with a fixed number of parameters, such that
$|f(x)-h(x)|<\varepsilon\quad\text{ for any }x\in[0,1].$ (8)
Figure 4: A sketch of our proof. In Steps 1 and 2, through interval partition, the approximation problem is transformed into a point-fitting problem. In Step 3, we use task-driven neurons to construct a unimodal function that can induce the dense trajectory. Then, composing the unimodal function can solve the point-fitting problem.
Proof sketch: Our proof leverages heavily the framework of [18] that turns via the interval partition the function approximation into the point-fitting problem. Hence, for consistency and readability, we inherit their notations and framework. The proof idea is divided into three steps:
Step 1. As Figure 4 shows, in equal distance, divide $[0,1)$ into small intervals $\mathcal{J}_{k}=\left[\frac{k-1}{K},\frac{k}{K}\right)$ for $k\in\{1,2,\cdots,K\}$, where $K$ is the number of intervals. The left endpoint of $\mathcal{J}_{k}$ is $x_{k}$. A higher $K$ leads to a smaller approximation error. We construct a piecewise constant function $h$ to approximate $f$. The error can be arbitrarily small as long as the divided interval goes tiny. Based on the interval division, we have
$h(x)\approx f(\tilde{x}_{k}),~{}~{}x\in\mathcal{J}_{k},$ (9)
where $\tilde{x}_{k}$ is a point from $\mathcal{J}_{k}$.
[…]
Step 3. In [18], the point-fitting function is constructed as $\tau(\theta/(\pi+n))$ based on Proposition 1. In contrast, we solve it with a discrete dynamic system that is constructed by a network made of task-driven neurons. We design a sub-network to generate a function $\phi_{2}$ mapping $k$ approximately to $f\left(\tilde{x}_{k}\right)$ for each $k$. Then $\phi_{2}(\phi_{1}(x))=\phi_{2}(k)\approx f(\tilde{x}_{k})\approx f(x)$ for any $x\in\mathcal{J}_{k}$ and $k\in\{1,2,\cdots,K\}$, which implies $\phi_{2}\circ\phi_{1}\approx f$ on $[0,1)$. $\phi_{2}$ is $T^{m(k)}(\xi)$ based on Lemma 1, where $T$ is a unimodal function from $[0,1]\to[0,1]$ that can also induce the dense trajectory.
Proposition 1 ([56]).
For any given set $\{w_{n}\}_{n=1}^{N}\subseteq[0,1]$ and $\epsilon>0$, there exists a value $\theta^{*}\in[0,+\infty)$, such that
$|w_{n}-\tau(\frac{\theta^{*}}{\pi+n})|<\epsilon,\;n=1,\cdots,N,$ (10)
where $\tau(z)=z-\lfloor z\rfloor$.
Definition 1 ([57]).
A map $T:[0,1]\rightarrow[0,1]$ is said to be unimodal if there exists a turning point $\eta\in[0,1]$ such that the map $f$ can be expressed as
$T(x)=\min\left\{T_{\mathrm{L}}(x),T_{\mathrm{R}}(x)\right\}=\left\{\begin{array}[]{l}T_{\mathrm{L}}(x),0\leqslant x\leqslant\eta,\\ T_{\mathrm{R}}(x),\eta\leqslant x\leqslant 1,\end{array}\right.$ (11)
where $T_{\mathrm{L}}:[0,\eta]\rightarrow[0,1]$ and $T_{\mathrm{R}}:[\eta,1]\rightarrow[0,1]$ are continuous, differentiable except possibly at finite points, monotonically increasing and decreasing, respectively, and onto the unit-interval in the sense that $T_{\mathrm{L}}(0)=T_{\mathrm{R}}(1)=0$ and $T_{\mathrm{L}}(\eta)=T_{\mathrm{R}}(\eta)=1$.
Lemma 1 (Dense Trajectory).
There exists a measure-preserving transformation $T$, generated from a network with task-driven neurons, such that there exists $\xi$ with $\{T^{n}(\xi):n=\mathbb{N}\}$ dense in $[0,1]$. In other words, for any $x$ and $\epsilon>0$, there exist $\xi$ and the corresponding composition times $m$, such that
$|x-T^{m}(\xi)|<\epsilon.$ (12)
Particularly, we refer to $\xi$ as a transitive point.
Proof.
Since we extensively apply polynomials to prototyping task-driven neurons, we prove that a polynomial of any order can fulfill the condition. The proof can be easily extended to other functions with minor twists.
i) Given a polynomial of any order, as shown in Figure 5, through cutting, translation, flipping, and scaling, we can construct a unimodal map $T(x):[0,1]\to[0,1]$.
Figure 5: A polynomial can be transformed into a unimodal map.
ii) For a unimodal map $T$, we can show that for any two non-empty sets $U,V\in X$ with $\mathfrak{m}(U),\mathfrak{m}(V)>0$, there exists an $m$ such that $T^{m}(U)\cap V\neq\varnothing$. As Figure 6 shows, this is because a set $U$ with $\mathfrak{m}(U)>0$ will always contain a tiny subset $\mathcal{J}_{mi}$ such that $T^{m}(\mathcal{J}_{mi})=[0,1]$. Then, naturally $T^{m}(U)\cap V\neq\varnothing$ due to $\mathcal{J}_{mi}\subset U$.
Figure 6: Given a set $U$ with $\mathfrak{m}(U)>0$, it will always contain a tiny subset $\mathcal{I}_{mk}$ such that $T^{mk}(U)=[0,1]$. Then, naturally $T^{m}(U)\cap V\neq\varnothing$, since $V\in[0,1]$.
iii) The above property can lead to a dense orbit, combining that $[0,1]$ is with a countable basis, according to Proposition 2 of [58]. For the sake of self-sufficiency, we include their proof here.
Let $\left(V_{i}\right)_{i\in I}$ be a countable base for $X$. For $i\in I$, the set $W_{i}=\bigcup_{n\geqq 0}f^{-n}\left(V_{i}\right)$ is open by continuity of $f$. This set is also dense in $X$. To see this let $U$ be any non-empty open set. Because of topological transitivity there exists a $k>0$ with $f^{k}(U)\cap V_{i}\neq\emptyset$. This gives $f^{-k}\left(V_{i}\right)\cap U\neq\emptyset$ and $W_{i}\cap U\neq\emptyset$. Thus $W_{i}$ is dense. By the Baire category theorem, the set $B=\cap_{i\in I}W_{i}$ is dense in $X$. Now, the orbit of any point $\xi\in B$ is dense in $X$.
Because, given any non-empty open $U\subset X$, there is an $i\in I$ with $V_{i}\subset U$ and $k>0$ with $\xi\in f^{-k}\left(V_{i}\right)$ This means $f^{k}(\xi)\in V_{i}\subset U$. Thus the orbit of $\xi$ enters any $U$.
Proof of Theorem 8. Now, we are ready to prove Theorem 8. First, the input $x$ is mapped into an integer $k$, whereas $k$ corresponds to the composition time $m(k)$. Next, a network module representing $T$ is copied $m(k)$ times and composed $m(k)$ times such that $T^{m(k)}(\xi)\approx f(\tilde{x}_{k})$. Since $f(\tilde{x}_{k})$ can approximate $f(x)$ well, $T^{m(k)}(\xi)$ also approximates $f(x)$ well.
Theorem 8 can be easily generalized into multivariate functions by using the Kolmogorov-Arnold theorem.
Proposition 2.
For any $\mathbf{x}=\left[x_{1},x_{2},\cdots,x_{d}\right]^{T}\in[0,1]^{d}$, there exist continuous functions $h_{i,j}\in C([0,1])$ for $i=0,1,\cdots,2d$ and $j=1,2,\cdots,d$ such that any continuous function $f\in C\left([0,1]^{d}\right)$ can be represented as
$f(\mathbf{x})=\sum_{i=0}^{2d}\Psi_{i}\left(\sum_{j=1}^{d}\Phi_{i,j}\left(x_{j}\right)\right).$ (13)
where $\Psi_{i}:\mathbb{R}\rightarrow\mathbb{R}$ is a continuous function for each $i\in\{0,1,\cdots,2d\}$.
Remark 2. Our construction saves the parametric complexity. In [18], though the number of parameters remains unchanged, the values of parameters $\theta^{*}$ increase as long as $n$ increases. Therefore, the entire parametric complexity goes up. In contrast, in our construction, the parameter value of $T$ is also fixed for different $(k,f(\tilde{x}_{k}))$. Yet, $T$ is copied and composed $m(k)$ times. The overall parametric complexity remains unchanged. What increases herein is computational complexity. Therefore, we call for brevity our construction “super-super-expressive”.
TABLE II: Regression results (MSE) on synthetic data. $\delta$ represents the noise level.
Ground Truth Formula $\delta=0\%$ $\delta=10\%$ $\delta=20\%$ Time (s)
SR TD SR TD SR TD SR TD
$2\mathbf{x}$ 0.0013 0.0008 0.0108 0.0120 0.0232 0.0138 13.94 4.46
$1.5(x\odot x)$ 0.0007 0.0006 0.0163 0.0095 0.0191 0.0166 12.75 6.69
$(\mathbf{x}\odot^{2}\mathbf{x})-1.5(\mathbf{x}\odot x)+2$ 0.0235 0.0028 0.2489 0.0102 0.1905 0.1706 28.77 7.73
$(\mathbf{x}\odot^{2}\mathbf{x})-1.5(\mathbf{x}\odot\mathbf{x})+3\mathbf{x}$ 0.0016 0.0012 0.0139 0.0101 0.0147 0.0116 22.01 8.82
$(\mathbf{x}\odot^{2}\mathbf{x})+(\mathbf{x}\odot\mathbf{x})$ 0.0095 0.0067 0.0116 0.0086 0.0495 0.0064 15.16 5.80
$4(\mathbf{x}\odot^{3}\mathbf{x})$ 0.0011 0.0002 0.0230 0.0165 0.0561 0.0193 24.77 9.44
$2(\mathbf{x}\odot^{3}\mathbf{x})+(\mathbf{x}\odot^{2}\mathbf{x})$ 0.0055 0.0014 0.0125 0.0104 0.0416 0.0109 15.80 8.21
$(\mathbf{x}\odot^{4}\mathbf{x})$ 0.0102 0.0062 0.0134 0.0111 0.0170 0.0126 13.30 8.24
Figure 7: Uniqueness experiment on airfoil self-noise data.
V Comparative Experiments on NeuronSeek
This section presents a comprehensive experimental evaluation of our proposed method and the previous NeuronSeek-SR [12]. First, we systematically demonstrate the superior performance and enhanced stability of NeuronSeek-TD compared to the previous NeuronSeek-SR. The results demonstrate that pursuing stability in NeuronSeek-TD does not hurt the representation power of the resultant task-driven neurons and even present better results.
V-A Experiment on Synthetic Data
Here, we adopt the synthetic data to compare the performance of NeuronSeek-TD and NeuronSeek-SR. The synthetic data are generated by adding noise perturbations to the known ground-truth functions. We then measure the regression error (MSE) of the formulas of NeuronSeek-SR and NeuronSeek-TD on the synthetic data.
V-A1 Experimental settings
Eight multivariate polynomials are constructed as shown in Table II. For simplicity and conciseness, we prescribe:
$\sum_{i=1}^{d}rx_{i}^{s}=r\cdot\underbrace{(\mathbf{x}\odot\cdots\odot\mathbf{x})^{\top}}_{s\text{ times}}=r\cdot(\mathbf{x}\odot^{s}\mathbf{x})^{\top},$
where $\bm{r}=[r,\ldots,r]\in\mathbb{R}^{d}$ denotes a constant vector, $\bm{x}=[x_{1},\ldots,x_{d}]$ represents the input vector, and $\odot$ denotes the Hadamard product. The elements of $\bm{x}$ are uniformly sampled from the interval $[-1,1]$. We set $d=30$ to simulate multi-dimensional inputs.
To evaluate the noise robustness of the proposed methods, we introduce three noise levels through
$\bar{y}=y+\delta\cdot\mathrm{RMS}(y)\cdot\bm{\eta},$
where $y$ denotes the noise-free polynomial output, $\bar{y}$ is the noisy observation, $\bm{\eta}\sim\mathcal{N}(\bm{0},\bm{I})$ is a standard normal random vector, $\delta$ controls the relative noise level, and $\mathrm{RMS}(y)$ represents the root mean square (RMS) of the data:
$\mathrm{RMS}(y)=\sqrt{\frac{1}{N}\sum_{i=1}^{N}y_{i}^{2}}.$
Here, $N=1,000$ specifies the number of data points. For NeuronSeek-SR, we configure a population size of 1,000 with the maximum of 50 generations. NeuronSeek-TD is optimized using Adam for 50 epochs. The noise levels are set to $\delta\in\{0\%,10\%,20\%\}$ respectively. We quantify the regression performance using the mean squared error (MSE) of the discovered neuronal formulae. Lower MSE values indicate better preservation of the underlying data characteristics.
Figure 8: The uniqueness experiment on phoneme data.
V-A2 Experimental results
The performance comparison on the synthetic data is presented in Table II. Our proposed NeuronSeek-TD demonstrates consistent superiority over NeuronSeek-SR across all generated data, with particularly notable improvements on the data derived from the function $(\mathbf{x}\odot^{2}\mathbf{x})-1.5(\mathbf{x}\odot\mathbf{x})+2$. This performance advantage remains robust under varying noise levels.
To assess computational efficiency, we measure the average execution time for both methods. NeuronSeek-TD achieves substantial computational savings, requiring less than half the execution time of the SR-based approach while maintaining comparable solution quality. These results collectively demonstrate that our method achieves both superior accuracy and greater computational efficiency.
V-B Experiments on Uniqueness and Stability
Here, we demonstrate that NeuronSeek-SR exhibits instability in terms of resulting in divergent formulas under minor perturbations in initialization, while NeuronSeek-TD maintains good consistency.
V-B1 Experimental settings
Our experimental protocol consists of two benchmark datasets: the phoneme dataset (a classification task classifying nasal and oral sounds) and the airfoil self-noise dataset (a regression task modeling self-generated noise of airfoil). These public datasets are chosen from the OpenML website to represent both classification and regression tasks, ensuring a comprehensive evaluation across different problem types. For each dataset, we introduce four levels of Gaussian noise ($\sigma\in\{0.01,0.025,0.05,0.1\}$) to simulate real-world perturbations. We train 10 independent instances of each method under identical initialization. For comparison, NeuronSeek-SR based on GP uses a population of 3,000 candidates evolved over 20 generations, whereas NeuronSeek-TD methods—employing CP decomposition and Tucker decomposition—rely on a model optimized via Adam over 20 epochs.
To quantify the uniqueness of SR and TD methods, we track two diversity metrics:
  • Epoch-wise diversity represents the number of different formulas within each epoch (or generation) across all initializations. For instance, $x^{3}+2x^{2}$ and $2x^{3}+x^{2}$ are considered identical, whereas $x^{3}+x^{2}$ and $x^{3}+x$ are distinct.
  • Cumulative diversity measures the average number of unique formulas identified per initialization as the optimization goes on.
V-B2 Experimental results
The results of phoneme and airfoil self-noise datasets are depicted in Figures 7 and 8, respectively. From the epoch-wise diversity (Figure 7 (a) and 8 (a)), we have several important observations: Firstly, GP-based SR demonstrates a rapid increase in the number of discovered formulas within each epoch. After training, it produces more unique formulas compared to TD-based methods (CP and Tucker). This indicates that SR is sensitive to initialization conditions. Secondly, as noise levels increase, the GP method discovers nearly 10 distinct formulas across 10 initializations. With higher noise, every initialization produces a completely different formula structure. These behaviors highlight the instability of GP-based SR when dealing with noisy data, as minor perturbations in initialization or data noise can lead to substantially different symbolic expressions.
TABLE III: Comparison of the network built by neurons using formulas generated by NeuronSeek-SR and NeuronSeek-TD. The number $h_{1}-h_{2}-\cdots-h_{k}$ means that this network has $k$ hidden layers, each with $h_{k}$ neurons. For example, 5-3-1 means this network has two hidden layers with 5 and 3 neurons respectively in each layer.
Datasets Instances Features Classes Metrics Structure Test results
TN NeuronSeek-TD NeuronSeek-SR NeuronSeek-TD
California housing 20640 8 continuous MSE 6-4-1 6-3-1 0.0720 $\pm$ 0.0024 0.0701 $\pm$ 0.0014
house sales 21613 15 continuous 6-4-1 6-4-1 0.0079 $\pm$ 0.0008 0.0075 $\pm$ 0.0004
airfoil self-noise 1503 5 continuous 4-1 3-1 0.0438 $\pm$ 0.0065 0.0397 $\pm$ 0.0072
diamonds 53940 9 continuous 4-1 4-1 0.0111 $\pm$ 0.0039 0.0102 $\pm$ 0.0015
abalone 4177 8 continuous 5-1 4-1 0.0239 $\pm$ 0.0024 0.0237 $\pm$ 0.0027
Bike Sharing Demand 17379 12 continuous 5-1 5-1 0.0176 $\pm$ 0.0026 0.0125 $\pm$ 0.0146
space ga 3107 6 continuous 2-1 2-1 0.0057 $\pm$ 0.0029 0.0056 $\pm$ 0.0030
Airlines DepDelay 8000 5 continuous 4-1 4-1 0.1645 $\pm$ 0.0055 0.1637 $\pm$ 0.0052
credit 16714 10 2 ACC 6-2 6-2 0.7441 $\pm$ 0.0092 0.7476 $\pm$ 0.0092
heloc 10000 22 2 18-2 17-2 0.7077 $\pm$ 0.0145 0.7089 $\pm$ 0.0113
electricity 38474 8 2 5-2 5-2 0.7862 $\pm$ 0.0075 0.7967 $\pm$ 0.0087
phoneme 3172 5 2 8-5 7-5 0.8242 $\pm$ 0.0207 0.8270 $\pm$ 0.0252
MagicTelescope 13376 10 2 6-2 6-2 0.8449 $\pm$ 0.0092 0.8454 $\pm$ 0.0079
vehicle 846 18 4 10-2 10-2 0.8176 $\pm$ 0.0362 0.8247 $\pm$ 0.0354
Oranges-vs-Grapefruit 10000 5 2 8-2 7-2 0.9305 $\pm$ 0.0160 0.9361 $\pm$ 0.0240
eye movements 7608 20 2 10-2 10-2 0.5849 $\pm$ 0.0145 0.5908 $\pm$ 0.0164
Furthermore, regarding cumulative diversity (Figures 7(b) and 8(b)), we can observe that GP-based SR constantly discovers an increasing number of different formulas, while TD methods tend to stabilize during training. This indicates that the GP algorithm struggles to converge to a consistent symbolic formula. In contrast, the asymptotic behavior exhibited by TD-based methods demonstrates their superior stability and convergence properties when identifying symbolic expressions. Moreover, the CP-based method suggests the best resilience against initialization variability.
V-C Superiority of NeuronSeek-TD over NeuronSeek-SR
To validate the superior performance of NeuronSeek-TD over NeuronSeek-SR, we conduct experiments across 16 benchmark datasets—8 for regression and 8 for classification—selected from scikit-learn and OpenML repositories. First, we use NeuronSeek-TD and NeuronSeek-SR simultaneously to search for the optimal formula. Second, we construct fully connected networks (FCNs) with neurons taking these discovered formulas as their aggregation functions. To verify the ability of NeuronSeek-TD, the structures of neural networks are carefully set, such that NeuronSeek-TD has equal or fewer parameters than NeuronSeek-SR.
Comprehensive implementation details and test results are provided in Table III. For regression tasks, the dataset is partitioned into training and test sets in an 8:2 ratio. The activation function is ReLU with MSE as the loss function and RMSProp as the optimizer. For classification tasks, the same train-test split is applied, using the Sigmoid activation function. The loss function and optimizer are set to CE and Adam, respectively. All experiments are repeated 10 times and reported by mean and standard deviation, with MSE for regression and accuracy for classification.
The regression results in Table III demonstrate that NeuronSeek-SR exhibits larger fitting errors compared to NeuronSeek-TD, along with higher standard deviations in MSE values. A notable example is the airfoil self-noise dataset, where NeuronSeek-TD outperforms NeuronSeek-SR by a significant margin, achieving a 9% reduction in MSE. Furthermore, NeuronSeek-TD shows superior performance on both the California housing and abalone datasets while utilizing fewer parameters. These results collectively indicate that NeuronSeek-TD possesses stronger regression capabilities than NeuronSeek-SR.
This performance advantage extends to classification tasks, where NeuronSeek-TD consistently achieves higher accuracy than NeuronSeek-SR. Particularly in the HELOC and orange-vs-grapefruit datasets, NeuronSeek-TD attains better classification performance with greater parameter efficiency. The observed improvements across both regression and classification tasks suggest that NeuronSeek-TD offers superior modeling capabilities compared to NeuronSeek-SR.
V-D Comparison of Tensor Decomposition Methods
Our framework employs the CP decomposition for neuronal discovery. To empirically validate this choice, we conduct a comparative analysis of some tensor decomposition methods used in NeuronSeek-TD.
V-D1 Experimental settings
Generally, we follow the previous experimental settings in Section V-C: 16 tabular datasets and the same network structures as described in Table III. We adopt CP decomposition and Tucker decomposition respectively for tensor processing in NeuronSeek-TD. We employ MSE for regression tasks, while classification accuracy is utilized for classification tasks.
Figure 9: Comparison between CP and Tucker decomposition. (a) Performance comparison on regression tasks measured by MSE. (b) Performance comparison on classification tasks measured by accuracy.
Figure 10: Comparison between CP and Tucker decomposition
V-D2 Experimental results
First, Figure 9 illustrates that, for most datasets, the two decomposition methods exhibit comparable performance. Exceptions include airfoil Self-Noise and electricity datasets, where the two methods show slight performance differences. This observation suggests that variations in TD methods do not substantially impact the final results. Second, as shown in Figure 10, the CP-based method achieves a 42% reduction in computational time. This result verifies that the CP method offers a computational advantage over the Tucker method.
VI Competitive Experiments of NeuronSeek-TD over Other Baselines
VI-A Experiment on Tabular Data
To further illustrate the superiority of NeuronSeek-TD, we compare it with other state-of-the-art models over 4 real-world tabular tasks: Electron Collision Prediction [^1] https://www.kaggle.com/datasets/fedesoriano/cern-electron-collision-data, Asteroid Prediction [^2] https://www.kaggle.com/datasets/basu369victor/prediction-of-asteroid-diameter, Heart Failure Detection [^3] https://www.kaggle.com/datasets/fedesoriano/heart-failure-prediction/data, and Stellar Classification [^4] https://www.kaggle.com/datasets/fedesoriano/stellar-classification-dataset-sdss17). We adopt 7 advanced machine learning methods, namely XGBoost [59], LightGBM [60], CatBoost [61], TabNet [62], TabTransformer [63], FT-Transformer [64], DANETs [65] and NeuronSeek-SR [12] as the baselines to compare with NeuronSeek-TD. Moreover, our proposed model and NeuronSeek-SR replace standard neurons with what are discovered by NeuronSeek-TD and NeuronSeek-SR, respectively.
All datasets are randomly split into training, validation, and test sets in a 0.8:0.1:0.1 ratio. For regression tasks (electron collision and asteroid prediction), we use the MSE as an evaluation metric, while for classification tasks with imbalanced labels (heart failure detection and stellar classification), we adopt the F1 score. Detailed experiment information is shown in Table IV.
TABLE IV: Details of real-world tabular datasets.
Dataset Task Data Size Features
Electron Collision Prediction Regression 99,915 16
Asteroid Prediction Regression 137,636 19
Heart Failure Detection Classification 1,190 11
Stellar Classification Classification 100,000 17
TABLE V: The test results of different models on real-world datasets.
Method electron collision (MSE) asteroid prediction (MSE) heartfailure (F1) Stellar Classification (F1)
XGBoost 0.0094 $\pm$ 0.0006 0.0646 $\pm$ 0.1031 0.8810 $\pm$ 0.02 0.9547 $\pm$ 0.002
LightGBM 0.0056 $\pm$ 0.0004 0.1391 $\pm$ 0.1676 0.8812 $\pm$ 0.01 0.9656 $\pm$ 0.002
CatBoost 0.0028 $\pm$ 0.0002 0.0817 $\pm$ 0.0846 0.8916 $\pm$ 0.01 0.9676 $\pm$ 0.002
TabNet 0.0040 $\pm$ 0.0006 0.0627 $\pm$ 0.0939 0.8501 $\pm$ 0.03 0.9269 $\pm$ 0.043
TabTransformer 0.0038 $\pm$ 0.0008 0.4219 $\pm$ 0.2776 0.8682 $\pm$ 0.02 0.9534 $\pm$ 0.002
FT-Transformer 0.0050 $\pm$ 0.0020 0.2136 $\pm$ 0.2189 0.8577 $\pm$ 0.02 0.9691 $\pm$ 0.002
DANETs 0.0076 $\pm$ 0.0009 0.1709 $\pm$ 0.1859 0.8948 $\pm$ 0.03 0.9681 $\pm$ 0.002
NeuronSeek-SR 0.0016 $\pm$ 0.0005 0.0513 $\pm$ 0.0551 0.8874 $\pm$ 0.05 0.9613 $\pm$ 0.002
NeuronSeek-TD (ours) 0.0011 $\pm$ 0.0003 0.0502 $\pm$ 0.0800 0.9023 $\pm$ 0.03 0.9714 $\pm$ 0.001
Results are compared in Table V. It can be seen that TabTransformer demonstrates competitive performance on the particle collision dataset. However, its effectiveness sharply declines on the asteroid prediction. In contrast, CatBoost and TabNet exhibit more consistent yet suboptimal results across four datasets. Lastly, the proposed NeuronSeek-TD method is the best performer, achieving the lowest error rates and highest F1 scores uniformly across all four benchmarks. Notably, NeuronSeek-TD significantly outperforms FT-Transformer and DANET. This significant margin underscores the efficacy of its task-adaptive architecture in recognizing feature interactions without domain-specific tuning.
VI-B Experiments on Images
To evaluate the performance enhancement of NeuronSeek-TD in computer vision tasks, we integrate it into several established convolutional neural network architectures: ResNet [3], DenseNet [66], SeResNet [67], and GoogLeNet [68]. Notably, we replace the first convolutional layer of each network block with a TN convolutional layer employing the discovered formula $x+\sin x$. It should be noted that the original NeuronSeek-SR was not designed for high-dimensional image tasks due to convergence difficulties. Our implementation focuses on NeuronSeek-TD and follows the standardized training protocol from a popular image classification repository [^5] https://github.com/weiaicunzai/pytorch-cifar100, incorporating several advanced training techniques: Learning rate scheduling via WarmUpLR and MultistepLR, regularization through Label Smoothing, and optimization using SGD. We use grid search to fine-tune the learning rate from 0.5 to 0.05. All models are trained for 200 epochs with a batch size of 128. We evaluate the results using classification accuracy averaged over 10 independent runs.
We adopt four standard image datasets to validate the model: CIFAR 10/100 [69], SVHN [70], and STL10 [71]. The training and validation sets are split with a ratio of 0.9 : 0.1. In particular, images from the STL10 dataset are resized to $32\times 32$ to fit the input of the model. The classification results are summarized in Table VI.
Our experiments demonstrate that task-driven neuronal layers consistently enhance the baseline performance across multiple datasets. Notably, NeuronSeek-augmented ResNet18 achieves a 2% accuracy improvement on both CIFAR 100 and STL10 datasets. While SeResNet exhibits median performance on CIFAR and SVHN datasets, its performance degrades on STL10; however, TN layers compensate for this degradation by boosting its accuracy by 10%. These results suggest that task-driven neuronal formula discovery can effectively unlock the latent potential of fundamental network architectures.
TABLE VI: Accuracy of compared methods on small-scale image datasets.
CIFAR10 CIFAR100 SVHN STL10
ResNet18 0.9419 0.7385 0.9638 0.6724
NeuronSeek-ResNet18 0.9464 0.7582 0.9655 0.7005
DenseNet121 0.9386 0.7504 0.9636 0.6216
NeuronSeek-DenseNet121 0.9399 0.7613 0.9696 0.6638
SeResNet101 0.9336 0.7382 0.9650 0.5583
NeuronSeek-SeResNet101 0.9385 0.7720 0.9685 0.6560
GoogleNet 0.9375 0.7378 0.9644 0.7241
NeuronSeek-GoogleNet 0.9400 0.7519 0.9658 0.7298
VII Discussions and Conclusions
In this paper, we have proposed a novel method in the framework of NeuronSeek to enhance the capabilities of a neural network. Compared to neural architecture search (NAS) [72], NeuronSeek provides a different approach. Rather than modifying the overall model structure, NeuronSeek optimizes solely at the neuronal level by searching for optimal aggregation formulas. Firstly, we have introduced tensor decomposition to prototype stable task-driven neurons. Secondly, we have provided theoretical guarantees demonstrating that these neurons can maintain the super-super-expressive property. At last, through extensive empirical evaluation, we show that our proposed framework consistently outperforms baseline models while achieving the state-of-the-art results across diverse datasets and tasks. These advancements establish task-driven neurons as an effective addition in the design of task-specific neural network components, offering both theoretical soundness and practical efficacy.

### v2

@@
 Abstract
-Drawing inspiration from our human brain that designs different neurons for different tasks, recent advances in deep learning have explored modifying a network’s neurons to develop so-called task-driven neurons.
-Prototyping task-driven neurons (referred to as NeuronSeek) employs symbolic regression (SR) to discover the optimal neuron formulation and construct a network from these optimized neurons.
-Along this direction, this work replaces symbolic regression with tensor decomposition (TD) to discover optimal neuronal formulations, offering enhanced stability and faster convergence.
-Furthermore, we establish theoretical guarantees that modifying the aggregation functions with common activation functions can empower a network with a fixed number of parameters to approximate any continuous function with an arbitrarily small error, providing a rigorous mathematical foundation for the NeuronSeek framework.
-Extensive empirical evaluations demonstrate that our NeuronSeek-TD framework not only achieves superior stability, but also is competitive relative to the state-of-the-art models across diverse benchmarks.
+Drawing inspiration from the human brain’s functional specialization, some recent advances in NeuroAI have ventured into task-driven neurons, which tailor neuronal formulations to specific tasks.
+However, existing prototypes rely heavily on Symbolic Regression (SR) via Genetic Programming (GP), which is plagued by the intrinsic weaknesses of the genetic programming such as the instability in searching for the optimal solution.
+To transcend these bottlenecks, we propose a paradigm shift from evolutionary heuristics to differentiable optimization.
+We propose NeuronSeek via Tensor Decomposition (NS-TD), a novel framework that reformulates the discrete search for neuronal structures into a continuous, low-rank tensor optimization problem.
+This transformation not only guarantees convergence stability but also significantly accelerates the discovery of better aggregation functions.
+Furthermore, we provide a rigorous theoretical foundation by proving that modifying aggregation functions—while retaining common activation functions—grants the network a “super-super-expressive” capability: approximating any continuous function with an arbitrarily small error using a fixed number and magnitude of parameters.
+Extensive empirical evaluations demonstrate that NS-TD establishes a new state-of-the-art, not only strictly dominating traditional SR-based methods in stability but also exhibiting competitive performance relative to state-of-the-art machine learning models.
 The code is available at https://github.com/HanyuPei22/NeuronSeek.
-Index Terms:
-Neuronal diversity, tensor decomposition, task-driven neuron, super-expressive
-I Introduction
+keywords:
+Neuronal diversity , Tensor decomposition , Task-driven neuron , Neural network
+journal: Pattern Recognition
+[…]
+[cuhk]organization=Frontier of Artificial Networks (FAN) Lab, Department of Data Science,
+City University of Hong Kong, city=Hong Kong, country=SAR of China
+[…]
+[polyu]organization=Department of Industrial and Systems Engineering, The Hong Kong Polytechnic University, city=Hong Kong, country=SAR of China
+[…]
+[riken]organization=RIKEN Center for Advanced Intelligence Project, city=Tokyo, country=Japan
+[…]
+[hust]organization=Center for Mathematical Sciences, and School of Mathematics and Statistics,
+Huazhong University of Science and Technology, city=Wuhan, country=China
+[…]
+[polyu2]organization=Department of Applied Mathematics, The Hong Kong Polytechnic University , city=Hong Kong, country=SAR of China
+{highlights}
+We propose a stable neuron discovery paradigm by differentiable tensor optimization.
+We prove that task-driven neurons possess “super-super-expressive” property with fixed number and magnitude of parameters.
+Our method achieves superior performance and enhanced stability compared to traditional symbolic regression methods and other state-of-the-art baselines across diverse benchmarks.
+1 Introduction
 Deep learning has achieved significant success in a wide range of fields [3, 35].
-Credits for these accomplishments largely go to the so-called scaling law, wherein fundamental computational units such as convolutions and attention mechanisms are designed and scaled per an optimized architecture.
+Credits for these accomplishments largely go to the development of scaling law, wherein basic computational units such as convolutions and attention mechanisms are designed and then scaled per a large-scale architecture.
+The resultant large network will embrace higher intelligence than human’s feature engineering given well-curated big data.
 Prominent examples include ResNet [22], Transformer [60], and Mamba [40].
 Recently, motivated by neuroscience’s paramount contributions to AI, an emerging field called NeuroAI has garnered significant attention, which draws upon the principles of biological circuits in the human brain to catalyze the next revolution in AI [65].
-The ambitious goal of NeuroAI is rooted in the belief that the human brain, as one of the most intelligent systems, inherently possesses the capacity to address complex challenges in AI development [7].
-Hence, it can always serve as a valuable source of inspiration for guiding practitioners, despite the correspondence between biological and artificial neural networks sometimes being implicit.
+The ambitious goal of NeuroAI is rooted in the belief that the human brain, as one of the most intelligent systems, inherently possesses the capacity to address complex challenges in AI development [16, 44].
+Hence, it can serve as a valuable source of inspiration for practitioners, despite the correspondence between biological and artificial neural networks sometimes is implicit.
 Following NeuroAI, let us analyze the scaling law through the lens of brain computation.
@@
 The latter is a natural consequence of stem cells’ directed programming in order to facilitate the task-specific information processing.
-Inspired by this observation, extensive research has shown that incorporating neuronal diversity into artificial neural networks can markedly enhance their capabilities [7].
+Inspired by this observation, extensive research has shown that incorporating neuronal diversity into artificial neural networks can markedly enhance their capabilities [16, 63, 62].
 For instance, researchers have transitioned beyond traditional inner-product neurons by integrating diverse nonlinear aggregation functions, such as quadratic neurons [63], polynomial neurons [9], and dendritic neurons [62].
-Such networks have achieved superior performance in tasks like image recognition and generation, underscoring the promise and practicality of innovating a neural network at the neuronal level.
+Such networks have achieved superior performance in tasks like image recognition and generation, underscoring the promise and practicality of innovating a neural network at the level of basic computational units.
 Along this direction, driven by the task-specific nature of neurons in our brain, recent research [17] introduced a systematic framework for task-driven neurons.
@@
 Therefore, it enables neurons to be tailored for specific tasks through a two-stage process.
-As shown in Figure 1, in the first stage, the vectorized symbolic regression (SR)—an extension of symbolic regression [13, 14] that enforces uniform operations for all variables—identifies the optimal homogeneous neuronal formula from data.
+Figure 1: Task-driven neuron based on the vectorized symbolic regression (NS-SR).
+As shown in Fig.
+1, in the first stage, the vectorized symbolic regression (SR)—an extension of symbolic regression [52, 2, 48] that enforces uniform operations for all variables—identifies the optimal homogeneous neuronal formula from data.
 Unlike traditional linear regression and symbolic regression, the vectorized SR searches for coefficients and homogeneous formulas at the same time, which can facilitate parallel computing.
@@
 More favorably, task-driven neurons can still be connected into a network to fully leverage the power of connectionism.
-Hereafter, we refer to task-driven neurons using symbolic regression as NeuronSeek-SR.
+Hereafter, we refer to task-driven neurons using symbolic regression as NS-SR.
 NeuronSeek considers in tandem neuronal importance and scale.
-From design to implementation, it systematically integrates three cornerstone AI paradigms: behaviorism (the genetic algorithm runs symbolic regression), symbolism (symbolic regression distills a formula from data), and connectionism.
-Figure 1: Task-driven neuron based on the vectorized symbolic regression (NeuronSeek-SR).
-Despite this initial success, the following two major challenges remain unresolved in prototyping task-driven neurons.
-  • Non-deterministic and unstable convergence.
-While symbolic regression employs modified genetic programming (GP) to discover polynomial expressions [15, 16], its tree-based evolutionary approach struggles to converge in high-dimensional space.
-The resulting formulas demonstrate acute sensitivity to certain hyperparameters, such as initialization conditions, population size, and mutation rate.
-This may confuse users about whether the identified formula is truly suitable for the task or due to the random perturbation, therefore hurting the legibility of task-driven neurons.
+From design to implementation, it systematically integrates three cornerstone AI paradigms: behaviorism (the genetic algorithm runs symbolic regression), symbolism (symbolic regression searches a formula from data), and connectionism (task-based neurons are connected into a network).
+Despite this initial success, the following two major challenges remain unresolved in prototyping task-driven neurons:
+  • 1.
+    Instability of Discrete Heuristic Search.
+Existing frameworks rely on Genetic Programming (GP) to navigate the vast space of mathematical expressions.
+This discrete, tree-based heuristic approach suffers from non-deterministic convergence, particularly in high-dimensional feature spaces.
+The discovered formulas are often fragile—highly sensitive to random seed initialization, population size, and mutation rates—rendering the resulting neurons inconsistent and difficult to reproduce.
+This trial-and-error nature fundamentally hurts practitioners’ trust on task-driven neurons.
+  • 2.
     Necessity of Task-driven Neurons.
@@
 Such a unique and desirable property allows a network to achieve a precise approximation without increasing structural complexity.
-Given the tremendous theoretical advantages of adjusting activation functions, it is necessary to address the following issue: Can the super-expressive property be achieved via revising the aggregation function while retaining common activation functions?
+Since task-based neurons revise aggregation functions instead of activation functions, it is necessary to figure out if the super-expressive property can be achieved via revising the aggregation function while retaining common activation functions?
 In this study, we address the above two issues satisfactorily.
-On one hand, in response to the instability of NeuronSeek-SR, we propose a tensor decomposition (TD) method referred to as NeuronSeek-TD.
+On the one hand, in response to the instability of NS-SR, we utilize a tensor decomposition (TD) method for neuronal discovery, referred to as NS-TD.
 This approach begins by assuming that the optimal representation of data is encapsulated by a high-order polynomial coupled with trigonometric functions.
 We construct the basic formulation and apply TD to optimize its low-rank structure and coefficients.
-Therefore, we reformulate the unstable formula search problem (NeuronSeek-SR) into a stable low-rank coefficient optimization task.
+Therefore, we reformulate the unstable formula search problem (NS-SR) into a stable low-rank coefficient optimization task.
 To enhance robustness, we introduce the sparsity regularization in the decomposition process, automatically eliminating insignificant terms while ensuring the framework to identify the optimal formula.
 Additionally, rank regularization is employed to derive the simplest possible representation.
-The improved stability of our method stems from two key factors: i) Unlike conventional symbolic regression which suffers from complex hyperparameter tuning in genetic programming [20], NeuronSeek-TD renders significantly fewer tunable parameters.
-ii) It has been proven that tensor decomposition with rank regularization tends to yield a unique solution [21], thereby effectively addressing the inconsistency issues inherent in symbolic regression methods.
+The improved stability of our method stems from two key factors: i) Unlike conventional symbolic regression which suffers from complex hyperparameter tuning in genetic programming [42], NS-TD renders significantly fewer tunable parameters.
+ii) It has been proven that TD with rank regularization tends to yield a unique solution [33], thereby effectively addressing the inconsistency issues inherent in symbolic regression methods.
 On the other hand, we close the theoretical deficit by showing that task-driven neurons that use common activation such as ReLU can also achieve the super-expressive property.
@@
 In summary, our contributions are threefold:
-  • We introduce tensor decomposition to prototype stable task-driven neurons by enhancing the stability of the process of finding a suitable formula for the given task.
-Our study is a major modification to NeuronSeek.
-  • We theoretically show that task-driven neurons with common activation functions also enjoy the “super-super-expressive” property, which puts the task-driven neuron on a solid theoretical foundation.
-The novelty of our construction lies in that it fixes both the number and values of parameters, while the previous construction only fixes the number of parameters.
-  • Systematic experiments on public datasets and real-world applications demonstrate that NeuronSeek-TD achieves not only superior performance but also stability compared to NeuronSeek-SR.
-Moreover, its performance is also competitive compared to other state-of-the-art baselines.
-II Related Works
-II-A Neuronal Diversity and Task-driven Neurons
+  • 1.
+    Paradigm Shift to Differentiable Search.
+Our work suggests a paradigm shift in neuronal structure discovery by transitioning from discrete evolutionary heuristics to a differentiable tensor optimization framework.
+Extensive evaluations demonstrate that this continuous optimization approach significantly surpasses traditional symbolic regression methods, achieving superior fitting capability and search efficiency while guaranteeing convergence stability.
+  • 2.
+    Theoretical Guarantee of “Super-Super-Expressiveness”.
+We provide a rigorous proof that task-driven neurons, even when using standard activation functions (e.g., ReLU), possess the “super-super-expressive” property.
+Distinct from prior constructions that only fix the number of parameters, our theory demonstrates that approximation capability holds even when both the number and magnitude of parameters are fixed, solidifying the theoretical foundation of the field.
+  • 3.
+    Superior Performance and Robustness.
+Systematic experiments on public datasets and real-world applications demonstrate that NS-TD achieves not only superior performance but also enhanced stability compared to NS-SR and traditional symbolic regression methods in the task of neuronal structure search.
+Moreover, its predictive performance remains competitive relative to other state-of-the-art baselines.
+2 Related Works
+2.1 Neuronal Diversity and Task-driven Neurons
 Many groundbreaking AI advancements were inspired by the computing of biological neural systems.
@@
 The human brain is renowned for its extensive array of neurons, which vary significantly in morphology and functionality.
-This neuronal diversity underpins the brain’s intelligent behavior  [23, 24].
+This neuronal diversity underpins the brain’s intelligent behavior [12].
 However, traditional artificial neural networks often prioritize the scaling law that designs a fundamental unit and replicates it across large architectures.
-This approach pays little attention to neuronal diversity.
-Recently, the idea of introducing neuronal diversity has naturally emerged, which mainly modifies the aggregation function instead of the activation function in constructing new neurons.
+This approach underestimates the importance of the diversity of basic computational units.
+Recently, the idea of introducing neuronal diversity emerged, which mainly modifies the aggregation function instead of the activation function in constructing new neurons.
 This operation is well-grounded: Once an activation function is monotonic, a neuron’s decision boundary is exclusively shaped by its aggregation function.
 In particular, it replaces the inner product of a neuron with a nonlinear function such as a polynomial.
-Table I summarizes the recently proposed non-linear neurons.
+Table 1 summarizes the recently proposed non-linear neurons.
 Notably, the complexity of neurons in [67, 27, 43] is of $\mathcal{O}(n^{2})$, which is much larger than the conventional neuron, while neurons in [20, 4, 63, 14] favorably enjoy the linear parametric complexity.
-TABLE I: A summary of the recently-proposed neurons.
+Table 1: A summary of the recently-proposed neurons.
 $\sigma(\cdot)$ is the nonlinear activation function.
@@
 [10, 7, 8] focus on designing polynomial network architectures, our work differs by concentrating on task-specific neuronal construction and more diverse functional forms such as trigonometric functions ($\sin(\mathbf{x})$, $\cos(\mathbf{x})$).
-II-B Symbolic Regression and Deep Learning
+2.2 Symbolic Regression and Deep Learning
 Symbolic regression (SR) is an algorithm that searches for formulas from data [52].
-Recent research in SR includes learning partial differential equations (PDEs) from data [34, 35], improving the validity of SR on high-dimensional data [36], and invertible SR [37].
+Recent research in SR includes learning partial differential equations (PDEs) from data [32], improving the validity of SR on high-dimensional data [50], and invertible SR [57].
 Besides, the Kolmogorov-Arnold Network (KAN) is an exemplary work that combines symbolic regression and deep learning [38].
@@
 Building upon this theorem, KAN is a neural network to exploit this decomposition by learning $g_{i}$ and $h_{i,j}$, which can be regarded as a special symbolic regression.
-Currently, this KAN-type line has been successfully used in a myraid of domains, such as medical image analysis [39] and scientific discovery [40].
+Currently, this KAN-type line has been successfully used in a myriad of domains, such as medical image analysis [36] and scientific discovery [58].
+Task-based neurons have fundamental differences from KAN in two aspects: i) Our work modifies what a single neuron computes.
+KAN modifies how the entire graph is connected and parameterized.
+KANs completely remove the concept of a ”neuron” with a fixed activation function.
+Instead, they place learnable univariate functions on the edges (weights), which sticks to a fixed topology.
+Thus, our work is of two stages, which is flexible in network structure design, which can fully leverage the power of connectionism.
+ii) Our work explicitly follows a ”one-for-one” philosophy.
+Users must re-run VSR for each new task/dataset to prototype a specific neuron.
+In contrast, KAN aims to be a universal approximator architecture; users do not redesign the KAN formula for a specific task.
+Lastly, it is important to note that the concept of task‑based neurons [17]—using symbolic regression to pre‑design a neuron’s aggregation function—was developed independently and earlier.
+[17] was finished and submitted in Nov 2023.
+Due to the subsequent grant review process, its dissemination was delayed.
+Hence, the appearance of KAN precedes our arXiv posting, but not our initial conception.
+We view KAN as an exciting parallel development that, together with our work, underscores the importance of symbolic computing in neural networks.
 Moreover, deep learning is also applied to strengthen the performance of SR algorithms.
@@
 SymbolicGPT, a Transformer-based language model for symbolic regression [59], is also shown to be a competent model compared with the existing models in terms of accuracy and run time.
-To summarize, symbolism is one of the earliest and most influential paradigms in the history of AI [44, 45].
+To summarize, symbolism is one of the most influential paradigms in the history of AI [47].
 It is rooted in the idea that intelligence can be modeled through the manipulation of symbols—abstract representations of objects, concepts, or relationships in the real world.
 This paradigm assumes that reasoning and problem-solving can be achieved through symbolic manipulation.
-Symbolism and connectionism are highly complementary [46, 47].
+Symbolism and connectionism are highly complementary  [55].
 Incorporating techniques in symbolism can effectively solve the intrinsic problems in connectionism such as efficiency, robustness, and forgetting.
-Our work synergizes symbolic regression and deep learning at the neuronal level, which is a novel practice in fusing symbolism and connectionism.
-II-C Super-expressive Activation Functions
+Our work synergizes symbolic is a novel practice in fusing symbolism and connectionism.
+2.3 Super-expressive Activation Functions
 These days, super-expressive activation functions gain lots of traction in deep learning approximation theory [51].
@@
 [66] uses the bit extract technique, while ours uses the chaotic theory.
+2.4 Task-driven Neuron Based on Vectorized Symbolic Regression
+Symbolic Regression (SR) aims to discover a mathematical expression that best describes a given dataset.
+Conventional SR methods typically treat each input variable independently, searching for heterogeneous formulas for different features.
+While flexible, this approach leads to an exponential growth in search space complexity as the input dimension increases.
+To mitigate this, the concept of Vectorized SR was introduced in the previous work NS-SR [17].
+The core idea is to impose a homogeneity constraint, enforcing that all input variables share the same underlying transformation rule.
+This regularization significantly reduces the search space complexity from exponential to linear with respect to the input dimension.
+As illustrated in Fig.
+1, frameworks like NS-SR typically follow a two-stage process: i) identifying the optimal homogeneous neuronal formula by fitting task-specific data; 2) implementing the derived formula as the neuron’s aggregation function.
+However, existing implementations of Vectorized SR predominantly rely on Genetic Programming (GP) as the search engine.
+While effective in low-dimensional settings, GP-based optimization faces critical bottlenecks in high-dimensional deep learning tasks.
+The discrete nature of evolutionary search operators, such as crossover and mutation, struggles to traverse the vast combinatorial space efficiently, often leading to premature convergence to local optima.
+Furthermore, GP-based methods are inherently sensitive to hyperparameters and initialization, resulting in unstable and inconsistent formula discoveries across different runs.
+These limitations necessitate a more stable, differentiable approach for discovering neuronal formulas, which motivates our proposed NS-TD framework.
+3 Task-driven Neuron Based on Tensor Decomposition (NS-TD)
+3.1 Overview of NS-TD
+To address the instability of GP-based methods, we propose NS-TD, a differentiable framework for neuronal formula discovery.
+As illustrated in Fig.
+2, distinct from the entangled search space in NS-SR, we mathematically decouple the search process into two parallel streams: a Polynomial Stream that explicitly models the non-linear contributions of individual features (e.g., polynomial powers), and an Interaction Stream dedicated to capturing cross-feature correlations of varying orders via tensor decomposition.
+This separation ensures that the model can efficiently capture dominant signals without being overwhelmed by the complexity of high-order interactions.
 Figure 2: The overall framework of the proposed method.
@@
 In the second stage, the neuronal formula is parameterized and integrated into various neural network backbones for task-specific applications.
-III Task-driven Neuron Based on Tensor Decomposition (NeuronSeek-TD)
-III-A NeuronSeek-SR
-In this section, we briefly illustrate how NeuronSeek-SR creates novel neurons.
-It follows a two-stage process, as illustrated in Figure 1.
-First, the vectorized SR identifies the optimal homogeneous neuronal formula by fitting task-specific data.
-Second, the derived formula is implemented as the neuron’s aggregation function with learnable coefficients.
-III-A1 Vectorized SR
-Conventional SR encodes a mathematical expression as a tree-structured directed graph and utilizes GP to navigate the solution space [15].
-The GP optimization process employs biologically-inspired evolutionary operators, including: i) crossover operations that exchange subtrees between candidate solutions to maintain population diversity, and ii) mutation operations that introduce random modifications to explore novel regions of the function space.
-While this approach offers notable flexibility for symbolic optimization, it faces significant challenges such as computational complexity, formula heterogeneity, and overfitting when applied to high-dimensional problems.
-All these together are not ideal for the task of prototyping new neuronal formulations.
-Unlike the conventional SR, which discovers heterogeneous formulas for each variable, the vectorized SR regularizes all input variables to learn the same formula [13, 14].
-This single change brings huge advantages: i) Lower complexity.
-Homogeneous formulas generated by the vectorized SR have significantly fewer parameters, scaling linearly with the input dimensionality.
-ii) Faster computation and minor overfitting risks.
-By limiting the search space, the vectorized SR reduces the risk of overfitting and also makes the regression process faster and more computationally efficient, especially for high-dimensional data.
-III-A2 Parameterization
-The learned formulas from the vectorized SR are parameterized, i.e., making the coefficients trainable.
-This allows neurons to fine-tune their behavior in the context of a network.
-While task-based neurons alone may have limited approximation power, connecting them into a neural network framework will further amplify the power of neurons, thereby enabling efficient optimization and improving performance in handling complex tasks.
-Despite the success of the vectorized SR, its built-in tree-based evolutionary algorithm still suffers from poor convergence in high-dimensional space.
-Consequently, the derived formulas exhibit significant sensitivity to key hyperparameters—including initialization conditions, population size, and mutation rate—often producing inconsistent solutions across different runs.
-This inherent instability in the NeuronSeek-SR framework stems from the intrinsic issues of genetic algorithms.
-III-B NeuronSeek-TD
-In this work, we propose NeuronSeek-TD, a faster and more stable alternative to conventional vectorized SR for neuronal formula discovery.
-We reasonably assert that the maximum required polynomial order for any given task is bounded by a constant $N$.
-By applying tensor decomposition to optimize coefficients, we reformulate the unstable formula search problem into a stable low-rank coefficient optimization task.
-Furthermore, we augment the polynomial structure through elementary functions, providing greater design flexibility and improved performance in high-dimensional pattern recognition tasks.
-The overall framework is illustrated in Figure 2, which comprises two stages: search for the neuronal formula and construct the task-driven neural network.
-III-B1 Neuronal formula discovery
-In the first stage, we adopt a polynomial model to identify the expressions that best fit the task-specific data.
-For consistent processing across modalities, both tabular and image inputs are flattened into a vector representation $\mathbf{z}\in\mathbb{R}^{d}$.
-Then, given the input data $\mathbf{z}$, the initial neuronal formula is defined as follows:
-$f(\mathbf{z})$ $=C_{1}\cdot\beta_{j}+C_{2}\cdot\mathbf{z}^{\top}\mathbf{W}^{[1]}+C_{3}\cdot\mathbf{z}^{\top}\mathbf{W}^{[2]}\mathbf{z}$ (2)
-$+C_{4}\cdot\mathbf{W}^{[3]}\times_{1}\mathbf{z}\times_{2}\mathbf{z}\times_{3}\mathbf{z}+\cdots+C_{N}\cdot\mathbf{W}^{[N]}\prod_{n=1}^{N}\times_{n}\mathbf{z}$
-$+D\cdot\sin(\mathbf{z}),$
-where $\mathbf{W}^{[n]}$ are weights and $C_{j}$ are coefficients. To put it in a more compact way, we have
-$f(\mathbf{z})$ $=\sum_{n=1}^{N}\left(C_{n}\cdot\mathcal{W}^{[n]}\prod_{m=1}^{n}\mathbf{z}\right)+D\cdot\sin(\mathbf{z}),$ (3)
-where the order $N$ and the rank are two hyperparameters to be set appropriately.
-The main idea of this method is to outer-product a series of rank-1 tensors to represent high-order tensors so that the computation cost is reduced significantly.
-The tensor decomposition process could be formulated as follows:
-$\mathbf{W}^{[n]}=\sum_{r=1}^{R}\mathbf{a}_{r}^{(1)}\circ\mathbf{a}_{r}^{(2)}\circ\dots\circ\mathbf{a}_{r}^{(n)},$ (4)
-where $R$ is decomposition rank, and $n$ is the order of the decomposed tensor $\mathbf{W}$.
-Here, we use COMP/PARAFAC (CP) decomposition to execute this process.
-CP decomposition has been widely adopted for modeling multilinear relationships in machine learning applications [51, 31].
-In contrast to Tucker decomposition, CP decomposition provides both computational and structural advantages for our neuronal formula discovery framework.
-Computationally, CP decomposition requires only $O(nRd)$ parameters for an $n$-th order tensor of size $d^{n}$, compared to Tucker’s $O(R^{n}+nRd)$ complexity, enabling efficient handling of high-order polynomial terms.
-An example of CP decomposition is illustrated in Figure 4.
-While other tensor decomposition approaches could be explored (e.g., tensor ring decomposition, hierarchical Tucker decomposition) [52, 53], the efficiency and structural benefits of CP decomposition have already aligned well with the requirements of our polynomial modeling framework.
+3.2 Neuronal Structure Searching
+In the first stage, we aim to discover the optimal mathematical structure for the neuron’s aggregation function.
+Given the input vector $\mathbf{z}\in\mathbb{R}^{d}$, the function $f(\mathbf{z})$ is defined as a combination of polynomial, interaction, and periodic terms, regulated by learnable gates.
+The general formulation is expressed as
+$f(\mathbf{z})$ $=\underbrace{\sum_{k=1}^{K}g_{k}^{(p)}\left(\mathbf{w}_{k}^{\top}\mathbf{z}^{\odot k}\right)}_{\text{Polynomial Stream}}+\underbrace{\sum_{m=2}^{M}g_{m}^{(i)}\left(\sum_{r=1}^{R}\prod_{j=1}^{m}(\mathbf{a}_{r,j}^{\top}\mathbf{z})\right)}_{\text{Interaction Stream}}+\underbrace{g^{(s)}\cdot D\sin(\mathbf{z})}_{\text{Periodic Term}},$ (2)
+where $g^{(\cdot)}\in\{0,1\}$ denotes the binary gate for term selection.
+The first component is the Polynomial Stream, where $\mathbf{w}_{k}^{\top}\mathbf{z}^{\odot k}$ models the independent $k$-th order polynomial effect, with $\mathbf{z}^{\odot k}$ denoting the element-wise power.
+The second component corresponds to the Interaction Stream, which utilizes CP decomposition to approximate the $m$-th order feature interactions efficiently.
+Finally, the periodic term $D\sin(\mathbf{z})$ is included to capture high-frequency patterns.
+3.2.1 Computation of Stream Components
+In this section, we detail the computational mechanisms underlying the polynomial and interaction streams defined above.
+The computation for the polynomial and periodic streams is straightforward, as they primarily involve element-wise operations.
+The polynomial term computes $\mathbf{z}^{\odot k}$ by raising each element of the input vector to the power of $k$ independently, while the periodic term applies the sine function to the input vector.
+These operations maintain a linear computational complexity with respect to the input dimension.
+The core computational challenge lies in the interaction stream, which models the high-order coupling between features.
+A naive implementation of an $m$-th order interaction would require a dense weight tensor $\mathcal{W}^{[m]}\in\mathbb{R}^{d\times\dots\times d}$ with $d^{m}$ parameters.
+To mitigate this exponential complexity, we employ the Canonical Polyadic (CP) decomposition.
+This method factorizes the high-dimensional weight tensor into a summation of rank-1 component tensors.
+Mathematically, the weight tensor $\mathcal{W}^{[m]}$ is approximated as
+$\mathcal{W}^{[m]}=\sum_{r=1}^{R}\mathbf{a}_{r}^{(1)}\circ\mathbf{a}_{r}^{(2)}\circ\dots\circ\mathbf{a}_{r}^{(m)},$ (3)
+where $R$ is the decomposition rank, $\circ$ denotes the vector outer product, and $\mathbf{a}_{r}^{(j)}\in\mathbb{R}^{d}$ represents the factor vector for the $j$-th mode of the $r$-th rank component.
+An illustration of a rank-$R$ CP decomposition for a third-order tensor is presented in Fig.
+3.
 Figure 3: A rank-R CP decomposition of a third-order tensor.
-Moreover, the incorporation of the trigonometric term $D\sin(\mathbf{z})$ serves a pivotal purpose.
-It extends the model’s capacity to capture periodic patterns and high-frequency components that intrinsically appear in many real-world datasets [54].
-This is particularly crucial as pure polynomial expansions exhibit limited spectral sensitivity.
-We then impose sparsity constraints through $L_{1}$ regularization on both the coefficients $C_{i}$ and weight tensors $\mathbf{W}^{[i]}$ to promote computational efficiency.
-Specifically, since each tensor $\mathbf{W}^{[i]}$ is decomposed using CP decomposition as $\mathbf{W}^{[i]}=\sum_{r=1}^{R}\mathbf{a}_{r}^{(1)}\circ\mathbf{a}_{r}^{(2)}\circ\dots\circ\mathbf{a}_{r}^{(i)}$, we apply sparsity regularization directly to the factor matrices.
-Thus, the complete optimization objective becomes
-$\underset{\{\mathbf{a}_{r}^{(j)}\},C_{i}}{\min}\sum_{k=1}^{K}\mathcal{L}(f(\mathbf{x}_{k}),y_{k})+\lambda_{1}\|\mathbf{C}\|_{1}+\lambda_{2}\sum_{i=1}^{N}\sum_{r=1}^{R}\sum_{j=1}^{i}\|\mathbf{a}_{r}^{(j)}\|_{1},$ (5)
-where $\mathcal{L}(\cdot)$ is the task-specific loss function, set to mean squared error (MSE) for regression and cross-entropy (CE) loss for classification.
-This structured sparsity approach encourages the discovery of interpretable and simplified polynomial formulas while maintaining computational efficiency through the low-rank tensor representation.
-To further enhance the compactness of the discovered neuronal formula beyond the $L_{1}$ regularization’s sparsity induction, we introduce a post-training thresholding mechanism for explicit term selection.
-While the $L_{1}$ regularization effectively drives many coefficients toward zero, empirical observations reveal that a non-negligible portion of coefficients $C_{i}$ retain small but non-zero values, which cumulatively increase model complexity without substantial predictive contribution.
-The thresholding mechanism is a statistically grounded pruning strategy: After model convergence, we compute the mean $\mu_{|C|}$ and the standard deviation $\sigma_{|C|}$ of the absolute values of all coefficients $\{|C_{i}|\}_{i=1}^{N}$, then construct the final neuronal formula by retaining only terms whose coefficients satisfy $|C_{i}|\geq\tau$, where $\tau=\mu_{|C|}-\sigma_{|C|}$.
-This can further increase the stability of identified formulas.
-Remark 1.
-Three key features of the proposed method warrant emphasis.
-i) The higher-order polynomial architecture systematically enables the learning of the polynomials.
-It is noticed that Eq.
-(2) is rather inclusive, e.g., it identifies multi-order variable interactions.
-This structural flexibility facilitates sophisticated nonlinear approximations, aligning the model more closely with the underlying data distribution.
-ii) The integration of trainable coefficients $\mathbf{C}$ with the $L_{1}$ regularization imposes selective sparsity constraints on the polynomial expansion.
-Specifically, the regularization term $\mathbf{C}$ induces sparsity by driving coefficients of less predictive terms toward zero, thereby automatically pruning nonessential higher-order components and yielding a parsimonious architecture.
-iii) The complementary $L_{1}$ regularization on the coefficients $\mathbf{W}$ enforces structured sparsity in the parameter space.
-This dual regularization strategy enhances feature selection while simultaneously improving computational efficiency—a critical advantage in high-dimensional contexts.
-Collectively, these design choices balance expressive capacity with compactness, ensuring robust performance without overparameterization.
-III-B2 Task-driven network construction
-In the second stage, the discovered neuronal formula is first parameterized into neuron-like structures.
-The selected polynomial and trigonometric terms are assigned trainable weights and integrated into neural network architectures.
-For tabular data processing, we employ fully connected neural networks.
-A polynomial with the maximum polynomial order of 3 is expressed as
-${y}=\sigma(\mathbf{W}_{1}^{\top}\mathbf{z}+\mathbf{W}_{2}^{\top}\mathbf{z}\odot\mathbf{z}+\mathbf{W}_{3}^{\top}\mathbf{z}\odot\mathbf{z}\odot\mathbf{z}+{b}),$ (6)
-with $\sigma$ being the ReLU activation function. If a specific order term is absent in the neuronal formula, its corresponding weight and bias parameters are initialized to zero.
-Moreover, we extend our framework to convolutional neurons for images.
-As mentioned, the trigonometric term is incorporated to model periodic patterns and high-frequency components inherent in high-dimensional images.
-Consequently, the complete neuronal formula is structured as follows:
-${y}=\sigma(\mathbf{W}_{1}^{\top}*\mathbf{z}+\mathbf{W}_{2}^{\top}*\mathbf{z}\odot\mathbf{z}+\mathbf{W}_{3}^{\top}*\mathbf{z}\odot\mathbf{z}\odot\mathbf{z}+\mathbf{W}_{4}*\sin(\mathbf{z})+{b}),$ (7)
-where * denotes the convolution operation.
-Finally, we introduce a simple initialization strategy to enhance training stability.
-Followed by [33], the weight matrix $\mathbf{W}_{1}$ is initialized as $\mathbf{W}_{1}\sim\mathcal{N}(0,\sigma^{2})$, where $\sigma=\sqrt{1/32(k_{n})}$, while all other weights are initialized to a tiny number such as 0.001.
-This strategy ensures that the network initially behaves like linear neurons and gradually learns higher-order terms during training.
-This strategy has been demonstrated the effectiveness in training quadratic neurons [55].
-IV Super-expressive Property Theory of Task-driven Neuron
+This decomposition provides significant structural and computational advantages.
+By substituting Eq.
+(3) into the interaction term, the computation transforms from a high-cost tensor contraction into a series of efficient vector inner products, as formulated in Eq.
+(2).
+Computationally, this reduces the parameter requirements from $O(d^{m})$ to $O(mRd)$.
+In contrast to other decomposition methods such as Tucker decomposition, which has a complexity of $O(R^{m}+mRd)$, CP decomposition eliminates the core tensor, thereby ensuring that the model complexity scales linearly with the input dimension $d$.
+This efficiency allows NS-TD to handle high-order interactions in high-dimensional spaces that are otherwise computationally intractable.
+3.2.2 Differentiable Structure Search
+To effectively identify the optimal structural configuration from the comprehensive formulation in Eq.
+(2), we employ a differentiable search strategy based on the $L_{0}$ regularization framework proposed by Louizos et al.
+[39].
+By assigning a stochastic binary gate to each candidate component in the polynomial, interaction, and periodic streams, we transform the discrete term selection problem into a continuous optimization task.
+We parameterize the activation probability of each gate using a learnable structural parameter.
+This formulation enables the gradients to backpropagate directly from the task loss to these parameters, allowing the optimizer to jointly tune the model weights (i.e., $\mathbf{w}$ and $\mathbf{a}$) and the gate probabilities.
+The global objective function balances predictive accuracy with structural parsimony:
+$\mathcal{J}=\sum_{(\mathbf{z},y)\in\mathcal{D}}\mathcal{L}_{task}(f(\mathbf{z}),y)+\lambda\sum_{g}\mathcal{R}(g),$ (4)
+where $\mathcal{L}_{task}$ is the task-specific loss (e.g., MSE or Cross-Entropy), and the second term aggregates the expected $L_{0}$ norm penalty $\mathcal{R}(g)$ over all structural gates defined in Eq.
+(2).
+Through this mechanism, the model automatically suppresses redundant polynomial orders or interaction ranks by driving their associated gate probabilities to zero, effectively performing feature selection during training.
+Crucially, to adapt this mechanism to the sensitivity of high-order polynomial and interaction learning, we implement a tailored two-phase optimization protocol.
+Direct application of sparsity regularization from the onset often leads to the premature pruning of complex terms (e.g., high-rank interactions) before they can capture relevant features.
+To mitigate this, we initially freeze the structural gates in an open state, focusing the optimization exclusively on training the model tensors to capture the dataset characterization.
+This warm-up phase ensures that all candidate terms, particularly those in the interaction stream utilizing CP decomposition, develop meaningful representations.
+In the subsequent phase, we unfreeze the gate parameters and introduce the sparsity regularization, allowing the model to prune terms based on their fully developed representational capacity.
+3.3 Task-driven Neural Network Construction
+Upon identifying the optimal sparse structure via the differentiable search in the first stage, we proceed to construct the task-driven neural network.
+This phase instantiates the discovered symbolic structure into a learnable neuronal architecture.
+Specifically, we extract the active index sets $\mathcal{K}$ and $\mathcal{M}$ for the retained polynomial and interaction terms, along with the inclusion status of the periodic term.
+Distinct from the search phase, here we assign fresh, independent trainable weights to these selected components to maximize representation learning during the final training.
+Mathematically, the output $y$ of the constructed neuron is defined as
+$y=\sigma\Bigg($ $\sum_{k\in\mathcal{K}}\mathbf{W}_{k}^{(p)}(\mathbf{z}^{\odot k})+\sum_{m\in\mathcal{M}}\mathbf{W}_{m}^{(i)}\left(\sum_{r=1}^{R}\prod_{j=1}^{m}(\mathbf{a}_{r,j}^{\top}\mathbf{z})\right)+\mathbf{W}^{(s)}\sin(\mathbf{z})+b\Bigg),$ (5)
+where $\sigma(\cdot)$ is the activation function (e.g., ReLU).
+$\mathbf{W}^{(\cdot)}$ represents the re-initialized trainable weights for the preserved terms.
+This formulation unifies the processing of tabular and visual data: for tabular inputs, $\mathbf{W}$ indicates scalar coefficients or weight matrices; for image inputs, operations are extended to convolutions.
+To ensure training stability for these higher-order architectures, we employ a hierarchical initialization strategy inspired by Chrysos et al.
+[8].
+Parameters corresponding to linear terms (i.e., $k=1$) are initialized using a normal distribution $\mathcal{N}(0,\sigma^{2})$ to preserve variance.
+In contrast, weights for higher-order polynomial terms and interaction terms are initialized with values close to zero (e.g., $10^{-3}$).
+This approach ensures that the network initially approximates a linear model, gradually integrating non-linear and interactive complexities during the training process.
+The efficacy of such initialization has been validated in prior studies on quadratic neurons [15].
+4 Super-expressive Property Theory of Task-driven Neurons
 In this section, we formally prove that task-driven neurons that modify the aggregation function can also achieve the super-expressive property.
@@
 $|f(x)-h(x)|<\varepsilon\quad\text{ for any }x\in[0,1].$ (6)
-Figure 4: A sketch of our proof.
-In Steps 1 and 2, through interval partition, the approximation problem is transformed into a point-fitting problem.
-In Step 3, we use task-driven neurons to construct a unimodal function that can induce the dense trajectory.
-Then, composing the unimodal function can solve the point-fitting problem.
 Proof sketch: Our proof leverages heavily the framework of [54] that turns via the interval partition the function approximation into the point-fitting problem.
@@
 Step 1.
-As Figure 4 shows, in equal distance, divide $[0,1)$ into small intervals $\mathcal{J}_{k}=\left[\frac{k-1}{K},\frac{k}{K}\right)$ for $k\in\{1,2,\cdots,K\}$, where $K$ is the number of intervals.
+As Fig.
+4 shows, in equal distance, divide $[0,1)$ into small intervals $\mathcal{J}_{k}=\left[\frac{k-1}{K},\frac{k}{K}\right)$ for $k\in\{1,2,\cdots,K\}$, where $K$ is the number of intervals.
 The left endpoint of $\mathcal{J}_{k}$ is $x_{k}$.
@@
 where $\tilde{x}_{k}$ is a point from $\mathcal{J}_{k}$.
-[…]
+Step 2.
+As Fig.
+4 shows, use the function $\phi_{1}(x)=\lfloor Kx\rfloor$ mapping all $x$ in the interval $\mathcal{J}_{k}$ to $k$, where $\lfloor\cdot\rfloor$ is the flooring function.
+There exists a one-to-one correspondence between $k$ and $f(\tilde{x}_{k})$.
+Thus, via dividing intervals and applying the floor function, the problem of approximation is simplified into a point-fitting problem.
+Therefore, we just need to construct a point-fitting function to map $k$ to $f(\tilde{x}_{k})$.
+Figure 4: A sketch of our proof.
+In Steps 1 and 2, through interval partition, the approximation problem is transformed into a point-fitting problem.
+In Step 3, we use task-driven neurons to construct a unimodal function that can induce the dense trajectory.
+Then, composing the unimodal function can solve the point-fitting problem.
 Step 3.
@@
 Then $\phi_{2}(\phi_{1}(x))=\phi_{2}(k)\approx f(\tilde{x}_{k})\approx f(x)$ for any $x\in\mathcal{J}_{k}$ and $k\in\{1,2,\cdots,K\}$, which implies $\phi_{2}\circ\phi_{1}\approx f$ on $[0,1)$.
-$\phi_{2}$ is $T^{m(k)}(\xi)$ based on Lemma 1, where $T$ is a unimodal function from $[0,1]\to[0,1]$ that can also induce the dense trajectory.
+$\phi_{2}$ is $T^{m(k)}(\xi)$ based on Lemma 1, where $T$ is a unimodal function from $0,1]\to[0,1]$ that can also induce the dense trajectory.
 Proposition 1 ([29]).
@@
 A map $T:[0,1]\rightarrow[0,1]$ is said to be unimodal if there exists a turning point $\eta\in[0,1]$ such that the map $f$ can be expressed as
-$T(x)=\min\left\{T_{\mathrm{L}}(x),T_{\mathrm{R}}(x)\right\}=\left\{\begin{array}[]{l}T_{\mathrm{L}}(x),0\leqslant x\leqslant\eta,\\ T_{\mathrm{R}}(x),\eta\leqslant x\leqslant 1,\end{array}\right.$ (11)
+$T(x)=\min\left\{T_{\mathrm{L}}(x),T_{\mathrm{R}}(x)\right\}=\left\{\begin{array}[]{l}T_{\mathrm{L}}(x),0\leqslant x\leqslant\eta,\\ T_{\mathrm{R}}(x),\eta\leqslant x\leqslant 1,\end{array}\right$.
+(9)
 where $T_{\mathrm{L}}:[0,\eta]\rightarrow[0,1]$ and $T_{\mathrm{R}}:[\eta,1]\rightarrow[0,1]$ are continuous, differentiable except possibly at finite points, monotonically increasing and decreasing, respectively, and onto the unit-interval in the sense that $T_{\mathrm{L}}(0)=T_{\mathrm{R}}(1)=0$ and $T_{\mathrm{L}}(\eta)=T_{\mathrm{R}}(\eta)=1$.
@@
 The proof can be easily extended to other functions with minor twists.
-i) Given a polynomial of any order, as shown in Figure 5, through cutting, translation, flipping, and scaling, we can construct a unimodal map $T(x):[0,1]\to[0,1]$.
 Figure 5: A polynomial can be transformed into a unimodal map.
+i) Given a polynomial of any order, as shown in Fig. 5, through cutting, translation, flipping, and scaling, we can construct a unimodal map $T(x):[0,1]\to[0,1]$.
 ii) For a unimodal map $T$, we can show that for any two non-empty sets $U,V\in X$ with $\mathfrak{m}(U),\mathfrak{m}(V)>0$, there exists an $m$ such that $T^{m}(U)\cap V\neq\varnothing$.
-As Figure 6 shows, this is because a set $U$ with $\mathfrak{m}(U)>0$ will always contain a tiny subset $\mathcal{J}_{mi}$ such that $T^{m}(\mathcal{J}_{mi})=[0,1]$.
+As Fig.
+6 shows, this is because a set $U$ with $\mathfrak{m}(U)>0$ will always contain a tiny subset $\mathcal{J}_{mi}$ such that $T^{m}(\mathcal{J}_{mi})=[0,1]$.
 Then, naturally $T^{m}(U)\cap V\neq\varnothing$ due to $\mathcal{J}_{mi}\subset U$.
@@
 Now, the orbit of any point $\xi\in B$ is dense in $X$.
-Because, given any non-empty open $U\subset X$, there is an $i\in I$ with $V_{i}\subset U$ and $k>0$ with $\xi\in f^{-k}\left(V_{i}\right)$ This means $f^{k}(\xi)\in V_{i}\subset U$.
+Because, given any non-empty open $U\subset X$, there is an $i\in I$ with $V_{i}\subset U$ and $k>0$ with $\xi\in f^{-k}\left(V_{i}\right)$ .
+This means $f^{k}(\xi)\in V_{i}\subset U$.
 Thus the orbit of $\xi$ enters any $U$.
@@
 Since $f(\tilde{x}_{k})$ can approximate $f(x)$ well, $T^{m(k)}(\xi)$ also approximates $f(x)$ well.
-Theorem 8 can be easily generalized into multivariate functions by using the Kolmogorov-Arnold theorem.
+Theorem 6 can be easily generalized into multivariate functions by using the Kolmogorov-Arnold theorem [38].
 Proposition 2.
@@
 Remark 2.
-Our construction saves the parametric complexity.
+Our construction reduces the parametric complexity.
 In [54], though the number of parameters remains unchanged, the values of parameters $\theta^{*}$ increase as long as $n$ increases.
-Therefore, the entire parametric complexity goes up.
+Therefore, the overall parametric complexity still goes up, as it consumes more digits.
 In contrast, in our construction, the parameter value of $T$ is also fixed for different $(k,f(\tilde{x}_{k}))$.
@@
 Therefore, we call for brevity our construction “super-super-expressive”.
-TABLE II: Regression results (MSE) on synthetic data. $\delta$ represents the noise level.
-Ground Truth Formula $\delta=0\%$ $\delta=10\%$ $\delta=20\%$ Time (s)
-SR TD SR TD SR TD SR TD
-$2\mathbf{x}$ 0.0013 0.0008 0.0108 0.0120 0.0232 0.0138 13.94 4.46
-$1.5(x\odot x)$ 0.0007 0.0006 0.0163 0.0095 0.0191 0.0166 12.75 6.69
-$(\mathbf{x}\odot^{2}\mathbf{x})-1.5(\mathbf{x}\odot x)+2$ 0.0235 0.0028 0.2489 0.0102 0.1905 0.1706 28.77 7.73
-$(\mathbf{x}\odot^{2}\mathbf{x})-1.5(\mathbf{x}\odot\mathbf{x})+3\mathbf{x}$ 0.0016 0.0012 0.0139 0.0101 0.0147 0.0116 22.01 8.82
-$(\mathbf{x}\odot^{2}\mathbf{x})+(\mathbf{x}\odot\mathbf{x})$ 0.0095 0.0067 0.0116 0.0086 0.0495 0.0064 15.16 5.80
-$4(\mathbf{x}\odot^{3}\mathbf{x})$ 0.0011 0.0002 0.0230 0.0165 0.0561 0.0193 24.77 9.44
-$2(\mathbf{x}\odot^{3}\mathbf{x})+(\mathbf{x}\odot^{2}\mathbf{x})$ 0.0055 0.0014 0.0125 0.0104 0.0416 0.0109 15.80 8.21
-$(\mathbf{x}\odot^{4}\mathbf{x})$ 0.0102 0.0062 0.0134 0.0111 0.0170 0.0126 13.30 8.24
-Figure 7: Uniqueness experiment on airfoil self-noise data.
-V Comparative Experiments on NeuronSeek
-This section presents a comprehensive experimental evaluation of our proposed method and the previous NeuronSeek-SR [12].
-First, we systematically demonstrate the superior performance and enhanced stability of NeuronSeek-TD compared to the previous NeuronSeek-SR.
-The results demonstrate that pursuing stability in NeuronSeek-TD does not hurt the representation power of the resultant task-driven neurons and even present better results.
-V-A Experiment on Synthetic Data
-Here, we adopt the synthetic data to compare the performance of NeuronSeek-TD and NeuronSeek-SR.
-The synthetic data are generated by adding noise perturbations to the known ground-truth functions.
-We then measure the regression error (MSE) of the formulas of NeuronSeek-SR and NeuronSeek-TD on the synthetic data.
-V-A1 Experimental settings
-Eight multivariate polynomials are constructed as shown in Table II. For simplicity and conciseness, we prescribe:
-$\sum_{i=1}^{d}rx_{i}^{s}=r\cdot\underbrace{(\mathbf{x}\odot\cdots\odot\mathbf{x})^{\top}}_{s\text{ times}}=r\cdot(\mathbf{x}\odot^{s}\mathbf{x})^{\top},$
-where $\bm{r}=[r,\ldots,r]\in\mathbb{R}^{d}$ denotes a constant vector, $\bm{x}=[x_{1},\ldots,x_{d}]$ represents the input vector, and $\odot$ denotes the Hadamard product.
-The elements of $\bm{x}$ are uniformly sampled from the interval $[-1,1]$.
-We set $d=30$ to simulate multi-dimensional inputs.
-To evaluate the noise robustness of the proposed methods, we introduce three noise levels through
-$\bar{y}=y+\delta\cdot\mathrm{RMS}(y)\cdot\bm{\eta},$
-where $y$ denotes the noise-free polynomial output, $\bar{y}$ is the noisy observation, $\bm{\eta}\sim\mathcal{N}(\bm{0},\bm{I})$ is a standard normal random vector, $\delta$ controls the relative noise level, and $\mathrm{RMS}(y)$ represents the root mean square (RMS) of the data:
-$\mathrm{RMS}(y)=\sqrt{\frac{1}{N}\sum_{i=1}^{N}y_{i}^{2}}.$
-Here, $N=1,000$ specifies the number of data points.
-For NeuronSeek-SR, we configure a population size of 1,000 with the maximum of 50 generations.
-NeuronSeek-TD is optimized using Adam for 50 epochs.
-The noise levels are set to $\delta\in\{0\%,10\%,20\%\}$ respectively.
-We quantify the regression performance using the mean squared error (MSE) of the discovered neuronal formulae.
-Lower MSE values indicate better preservation of the underlying data characteristics.
-Figure 8: The uniqueness experiment on phoneme data.
-V-A2 Experimental results
-The performance comparison on the synthetic data is presented in Table II.
-Our proposed NeuronSeek-TD demonstrates consistent superiority over NeuronSeek-SR across all generated data, with particularly notable improvements on the data derived from the function $(\mathbf{x}\odot^{2}\mathbf{x})-1.5(\mathbf{x}\odot\mathbf{x})+2$.
-This performance advantage remains robust under varying noise levels.
-To assess computational efficiency, we measure the average execution time for both methods.
-NeuronSeek-TD achieves substantial computational savings, requiring less than half the execution time of the SR-based approach while maintaining comparable solution quality.
-These results collectively demonstrate that our method achieves both superior accuracy and greater computational efficiency.
-V-B Experiments on Uniqueness and Stability
-Here, we demonstrate that NeuronSeek-SR exhibits instability in terms of resulting in divergent formulas under minor perturbations in initialization, while NeuronSeek-TD maintains good consistency.
-V-B1 Experimental settings
+5 Analysis Experiments of NeuronSeek
+5.1 Experiment on Synthetic Data
+In this section, we compare the performance of our proposed NS-TD with former NS-SR[17] and traditional symbolic regression methods on synthetic data.
+We select representative benchmark methods including classic symbolic regression [52], EQL [45], and the state-of-the-art method MetaSymNet [37].
+Synthetic dataset is generated by the functions in Table 2.
+For each task, input vectors $\mathbf{x}\in\mathbb{R}^{d}$ are sampled from a standard normal distribution $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$, and the target $\mathbf{y}$ is standardized to zero mean and unit variance.
+Crucially, regarding the function coefficients, while most are sampled from a normal distribution to ensure diversity, we deliberately assign high-contrast values (e.g., 10 vs.
+0.5) in specific cases.
+This design simulates signal masking scenarios, thereby increasing the difficulty of capturing weak structural information amidst dominant features.
+Table 2: Synthetic Dataset Generation Formulas. Comparisons of Pure, Interaction, and Hybrid modes. Coefficients $c_{i}$ are sampled from $\mathcal{N}(0,1)$.
+ID Pure Mode Interact Mode Hybrid Mode
+$f(\mathbf{x})\sim P_{k}$ $f(\mathbf{x})\sim I_{k}$ $f(\mathbf{x})\sim P_{k}+I_{k}$
+0 $c\cdot P_{2}(\mathbf{x})$ $c\cdot I_{2}(\mathbf{x})$ $c_{1}P_{2}(\mathbf{x})+c_{2}I_{2}(\mathbf{x})$
+1 $c\cdot P_{3}(\mathbf{x})$ $c_{1}I_{2}(\mathbf{x})+c_{2}I_{3}(\mathbf{x})$ $c_{1}P_{1}(\mathbf{x})+c_{2}I_{2}(\mathbf{x})$
+2 $10P_{1}(\mathbf{x})+0.5P_{2}(\mathbf{x})$ $8I_{2}(\mathbf{x})+0.5I_{3}(\mathbf{x})$ $5P_{3}(\mathbf{x})+0.5I_{2}(\mathbf{x})$
+3 $5P_{2}(\mathbf{x})+0.5P_{4}(\mathbf{x})$ $c_{1}I_{2}(\mathbf{x})+c_{2}I_{4}(\mathbf{x})$ $0.5P_{2}(\mathbf{x})+5I_{3}(\mathbf{x})$
+4 $c\cdot P_{5}(\mathbf{x})$ $c\cdot I_{4}(\mathbf{x})$ $c_{1}P_{2}(\mathbf{x})+c_{2}I_{2}(\mathbf{x})+c_{3}I_{3}(\mathbf{x})$
+Note: $P_{k}(\mathbf{x})=\sum_{j=1}^{d}x_{j}^{k}$ denotes the element-wise pure polynomial term.
+$I_{k}(\mathbf{x})=\prod_{m=1}^{k}(\mathbf{w}_{m}^{\top}\mathbf{x})$ denotes the interaction term of order $k$, where $\mathbf{w}_{m}\sim\mathcal{N}(0,1/\sqrt{d})$.
+The experimental pipeline is structured into two distinct stages: structure discovery and structure evaluation.
+In the first stage, all methods operate under a strict time budget of 60 seconds to search for the underlying regression formula of a given dataset.
+Subsequently, we extract the symbolic structure of the discovered formula and proceed to the second stage, where we re-optimize the constants from scratch to evaluate the model’s representational capacity.
+This two stage method is designed to ensure that we solely evaluate the capacity of searched structure.
+For our proposed NS-TD, we set the tensor rank to 8, which provides sufficient capacity for the synthetic functions, relying on the model’s sparsity mechanism to prune redundant components.
+The regularization coefficient is set to 0.05, a value empirically selected to balance reconstruction fidelity with structural sparsity.
+For genetic programming-based methods (Standard-SR and TN-SR), following the configuration in [17], we set the population size to 1,000, with crossover, mutation, and reproduction probabilities of 30%, 60%, and 10%, respectively; an elitism strategy is applied to select the top 5% of individuals for the next generation.
+For MetaSymNet [37], we adopt an entropy loss regularization of 0.2 and a learning rate of 0.01.
+The hyperparameter settings for EQL follow those in [45].
+For data preparation, we generate 2,500 samples per function, split into training and test sets with an 80/20 ratio.
+We report the average MSE across five distinct functions in Table 3.
+NS-TD demonstrates a decisive advantage over competing symbolic regression baselines.
+Across all functional modes and dimensionalities, our method consistently achieves the lowest test MSE, validating its superior capability in identifying mathematical structures that accurately approximate high-dimensional data manifolds.
+In contrast, while the tensor network-based NS-SR outperforms traditional discrete search methods like PySR and MetaSymNet in the Pure Mode, it exhibits a significant performance deterioration in the Interact and Hybrid modes.
+This performance gap stems from NS-SR’s architectural reliance on univariate transformations, which precludes the modeling of cross-feature interactions, thereby failing to capture the coupled dynamics inherent in the latter two scenarios.
+Moreover, NS-TD displays remarkable stability as the input dimension scales from d=10 to d=100.
+Unlike baseline methods which suffer from convergence difficulties, our approach maintains consistently low error rates, confirming its scalability and robustness in handling structure discovery tasks with high-dimensional data, which is common in deep learning scenarios.
+Moreover, the Standard Error of the Mean (SEM) across different trials are reported in Table 911 in the Appendix.
+Table 3: Performance Analysis on Synthetic Data.
+We evaluate the reconstruction accuracy (Mean MSE) across Pure, Interact, and Hybrid modes under varying input dimensions ($d\in\{10,30,50,100\}$).
+Best results are highlighted in bold.
+Method Pure Mode Interact Mode Hybrid Mode
+($\sum x^{k}$) (Vector Interact) (Mixture)
+10 30 50 100 10 30 50 100 10 30 50 100
+SR (PySR) 1.1612 1.1493 1.2325 1.2481 0.8754 1.0042 0.9776 0.9763 1.0635 0.9821 1.0764 1.1018
+EQL 0.4524 0.5342 0.5705 0.6923 0.3875 0.4862 0.5614 0.4521 0.3126 0.4113 0.4237 0.4472
+NS-SR (VSR) 0.2324 0.2541 0.2825 0.2953 0.6972 0.7734 0.7326 0.8145 0.7783 0.7824 0.8281 0.8685
+MetaSymNet 0.8324 0.8541 0.9825 0.9843 0.8632 1.0804 1.0193 0.9942 1.1345 0.9526 1.1093 1.1334
+NS-TD 0.0543 0.0632 0.0854 0.0891 0.0512 0.0583 0.0724 0.0945 0.0423 0.0621 0.0735 0.1112
+Convergence and Efficiency.
+To further investigate the optimization dynamics, we analyze the convergence trajectory of all methods on the most challenging task (Hybrid Mode, $d=100$), as illustrated in Fig.
+7.
+We can observe that our method NS-TD (Red line) converges to the optimal solution faster than evolutionary baselines.
+This confirms that our differentiable architecture search effectively bypasses the combinatorial search space that hinders methods like NS-SR and SR.
+Though EQL stabilizes at a plateau at 5.5 seconds, quicker than NS-TD (approximately at 10s), this higher plateau indicates that EQL is stuck in local optima.
+Figure 7: Convergence Efficiency Analysis (Hybrid Mode, $d=100$).
+Why Traditional Symbolic Regression Fails?
+The failure of baseline methods stems from two fundamental limitations.
+First, the combinatorial explosion in high-dimensional search spaces prevents discrete search algorithms from converging into a good solution within practical time budgets.
+Second, and more critically, traditional SR tends to generate structurally irregular formulas that are mathematically incompatible with neural architectures that favor homogeneous formulas.
+Due to the prohibitive length and structural heterogeneity of baseline outputs, we present a representative case study in Table 4.
+As illustrated, baseline methods suffer from consistently observed structural pathologies across high-dimensional tasks: SR generates index-dependent heterogeneous terms; EQL results in excessive density; and MetaSymNet collapses into local features.
+These heterogeneous structures violate the homogeneity required for tensor operations, meaning they cannot be efficiently vectorized or broadcasted across features.
+Consequently, unlike the compact forms discovered by NS-TD, these irregular baselines are computationally intractable to serve as scalable aggregation functions for dense deep learning layers.
+Deep learning tasks predominantly involve high-dimensional data representations.
+Through our experiments simulating these high-dimensional generation processes, NS-TD has demonstrated exceptional efficiency in identifying formulas that accurately capture the underlying data patterns.
+Crucially, in contrast to traditional symbolic regression approaches that yield irregular and computationally intractable expressions, the formulas discovered by NS-TD exhibit the structural homogeneity and compactness required for effective neuronal aggregation.
+This confirms that NS-TD uniquely satisfies the dual objectives of high reconstruction fidelity and structural suitability for constructing task-based neurons.
+Table 4: Case Study: Discovered Structures on Hybrid Mode (Formula ID:0, $d=100$).
+We illustrate the symbolic forms identified by different methods.
+While NS-TD successfully recovers the homogeneous vectorized structure, baselines generate heterogeneous or dense formulas that are ill-suited for neuronal aggregation.
+Method Discovered Structure
+Ground Truth $P_{2}(\mathbf{x})+I_{2}(\mathbf{x})$
+NS-TD (Ours) $P_{2}(\mathbf{x})+I_{2}(\mathbf{x})$
+NS-SR $P_{1}(\mathbf{x})+P_{2}(\mathbf{x})$
+SR $\frac{x_{12}+x_{19}}{x_{16}}+x_{25}^{2}-0.05x_{12}\dots$
+EQL Mixture of 79 active terms ($\sum w_{i}\phi_{i}(\mathbf{x})$)
+MetaSymNet $\log(\exp(x_{82}))+x_{45}^{2}$
+5.2 Experiments on Uniqueness and Stability
+Here, we demonstrate that NS-SR exhibits instability in terms of resulting in divergent formulas under minor perturbations in initialization, while NS-TD maintains good consistency.
+5.2.1 Experimental Settings
 Our experimental protocol consists of two benchmark datasets: the phoneme dataset (a classification task classifying nasal and oral sounds) and the airfoil self-noise dataset (a regression task modeling self-generated noise of airfoil).
@@
 We train 10 independent instances of each method under identical initialization.
-For comparison, NeuronSeek-SR based on GP uses a population of 3,000 candidates evolved over 20 generations, whereas NeuronSeek-TD methods—employing CP decomposition and Tucker decomposition—rely on a model optimized via Adam over 20 epochs.
+For comparison, NS-SR based on GP uses a population of 3,000 candidates evolved over 20 generations, whereas NS-TD methods—employing CP decomposition and Tucker decomposition—rely on a model optimized via Adam over 20 epochs.
 To quantify the uniqueness of SR and TD methods, we track two diversity metrics:
+  • 1.
     Epoch-wise Diversity represents the number of different formulas within each epoch (or generation) across all initializations.
 For instance, $x^{3}+2x^{2}$ and $2x^{3}+x^{2}$ are considered identical, whereas $x^{3}+x^{2}$ and $x^{3}+x$ are distinct.
+  • 2.
     Cumulative Diversity measures the average number of unique formulas identified per initialization as the optimization goes on.
-V-B2 Experimental results
-The results of phoneme and airfoil self-noise datasets are depicted in Figures 7 and 8, respectively.
-From the epoch-wise diversity (Figure 7 (a) and 8 (a)), we have several important observations: Firstly, GP-based SR demonstrates a rapid increase in the number of discovered formulas within each epoch.
+5.2.2 Experimental Results
+The results of phoneme and airfoil self-noise datasets are depicted in Figs. 8 and 9, respectively.
+(a) Epoch-wise diversity
+(b) Cumulative diversity (average per initialization)
+Figure 8: Uniqueness experiment on airfoil self-noise data.
+(a) Epoch-wise diversity
+(b) Cumulative diversity (average per initialization)
+Figure 9: Uniqueness experiment on phoneme data.
+From the epoch-wise diversity (Fig.
+8 (a) and Fig.
+9 (a)), we have several highlights: Firstly, GP-based SR demonstrates a rapid increase in the number of discovered formulas within each epoch.
 After training, it produces more unique formulas compared to TD-based methods (CP and Tucker).
@@
 These behaviors highlight the instability of GP-based SR when dealing with noisy data, as minor perturbations in initialization or data noise can lead to substantially different symbolic expressions.
-TABLE III: Comparison of the network built by neurons using formulas generated by NeuronSeek-SR and NeuronSeek-TD.
-The number $h_{1}-h_{2}-\cdots-h_{k}$ means that this network has $k$ hidden layers, each with $h_{k}$ neurons.
-For example, 5-3-1 means this network has two hidden layers with 5 and 3 neurons respectively in each layer.
-Datasets Instances Features Classes Metrics Structure Test results
-TN NeuronSeek-TD NeuronSeek-SR NeuronSeek-TD
-California housing 20640 8 continuous MSE 6-4-1 6-3-1 0.0720 $\pm$ 0.0024 0.0701 $\pm$ 0.0014
-house sales 21613 15 continuous 6-4-1 6-4-1 0.0079 $\pm$ 0.0008 0.0075 $\pm$ 0.0004
-airfoil self-noise 1503 5 continuous 4-1 3-1 0.0438 $\pm$ 0.0065 0.0397 $\pm$ 0.0072
-diamonds 53940 9 continuous 4-1 4-1 0.0111 $\pm$ 0.0039 0.0102 $\pm$ 0.0015
-abalone 4177 8 continuous 5-1 4-1 0.0239 $\pm$ 0.0024 0.0237 $\pm$ 0.0027
-Bike Sharing Demand 17379 12 continuous 5-1 5-1 0.0176 $\pm$ 0.0026 0.0125 $\pm$ 0.0146
-space ga 3107 6 continuous 2-1 2-1 0.0057 $\pm$ 0.0029 0.0056 $\pm$ 0.0030
-Airlines DepDelay 8000 5 continuous 4-1 4-1 0.1645 $\pm$ 0.0055 0.1637 $\pm$ 0.0052
-credit 16714 10 2 ACC 6-2 6-2 0.7441 $\pm$ 0.0092 0.7476 $\pm$ 0.0092
-heloc 10000 22 2 18-2 17-2 0.7077 $\pm$ 0.0145 0.7089 $\pm$ 0.0113
-electricity 38474 8 2 5-2 5-2 0.7862 $\pm$ 0.0075 0.7967 $\pm$ 0.0087
-phoneme 3172 5 2 8-5 7-5 0.8242 $\pm$ 0.0207 0.8270 $\pm$ 0.0252
-MagicTelescope 13376 10 2 6-2 6-2 0.8449 $\pm$ 0.0092 0.8454 $\pm$ 0.0079
-vehicle 846 18 4 10-2 10-2 0.8176 $\pm$ 0.0362 0.8247 $\pm$ 0.0354
-Oranges-vs-Grapefruit 10000 5 2 8-2 7-2 0.9305 $\pm$ 0.0160 0.9361 $\pm$ 0.0240
-eye movements 7608 20 2 10-2 10-2 0.5849 $\pm$ 0.0145 0.5908 $\pm$ 0.0164
-Furthermore, regarding cumulative diversity (Figures 7(b) and 8(b)), we can observe that GP-based SR constantly discovers an increasing number of different formulas, while TD methods tend to stabilize during training.
+Furthermore, regarding cumulative diversity (Fig.
+8(b) and Fig.
+9(b)), we can observe that GP-based SR constantly discovers an increasing number of different formulas, while TD methods tend to stabilize during training.
 This indicates that the GP algorithm struggles to converge to a consistent symbolic formula.
@@
 Moreover, the CP-based method suggests the best resilience against initialization variability.
-V-C Superiority of NeuronSeek-TD over NeuronSeek-SR
-To validate the superior performance of NeuronSeek-TD over NeuronSeek-SR, we conduct experiments across 16 benchmark datasets—8 for regression and 8 for classification—selected from scikit-learn and OpenML repositories.
-First, we use NeuronSeek-TD and NeuronSeek-SR simultaneously to search for the optimal formula.
+5.3 Superiority of NS-TD
+To validate the superior performance of NS-TD, we conduct experiments across 16 benchmark datasets—8 for regression and 8 for classification—selected from scikit-learn and OpenML repositories.
+This experiment is also conducted in 2 stages.
+First, we use NS-TD and NS-SR simultaneously to search for the optimal formula.
 Second, we construct fully connected networks (FCNs) with neurons taking these discovered formulas as their aggregation functions.
-To verify the ability of NeuronSeek-TD, the structures of neural networks are carefully set, such that NeuronSeek-TD has equal or fewer parameters than NeuronSeek-SR.
-Comprehensive implementation details and test results are provided in Table III.
+We evaluate the MSE on regression tasks and accuracy on classification tasks.
+For comparison, we select Kolmogorov-Arnold Network(KAN) [38] and multi-layer perceptron(MLP) as baselines.
 For regression tasks, the dataset is partitioned into training and test sets in an 8:2 ratio.
@@
 The loss function and optimizer are set to CE and Adam, respectively.
-All experiments are repeated 10 times and reported by mean and standard deviation, with MSE for regression and accuracy for classification.
-The regression results in Table III demonstrate that NeuronSeek-SR exhibits larger fitting errors compared to NeuronSeek-TD, along with higher standard deviations in MSE values.
-A notable example is the airfoil self-noise dataset, where NeuronSeek-TD outperforms NeuronSeek-SR by a significant margin, achieving a 9% reduction in MSE.
-Furthermore, NeuronSeek-TD shows superior performance on both the California housing and abalone datasets while utilizing fewer parameters.
-These results collectively indicate that NeuronSeek-TD possesses stronger regression capabilities than NeuronSeek-SR.
-This performance advantage extends to classification tasks, where NeuronSeek-TD consistently achieves higher accuracy than NeuronSeek-SR.
-Particularly in the HELOC and orange-vs-grapefruit datasets, NeuronSeek-TD attains better classification performance with greater parameter efficiency.
-The observed improvements across both regression and classification tasks suggest that NeuronSeek-TD offers superior modeling capabilities compared to NeuronSeek-SR.
-V-D Comparison of Tensor Decomposition Methods
-Our framework employs the CP decomposition for neuronal discovery. To empirically validate this choice, we conduct a comparative analysis of some tensor decomposition methods used in NeuronSeek-TD.
-V-D1 Experimental settings
-Generally, we follow the previous experimental settings in Section V-C: 16 tabular datasets and the same network structures as described in Table III.
-We adopt CP decomposition and Tucker decomposition respectively for tensor processing in NeuronSeek-TD.
-We employ MSE for regression tasks, while classification accuracy is utilized for classification tasks.
-Figure 9: Comparison between CP and Tucker decomposition. (a) Performance comparison on regression tasks measured by MSE. (b) Performance comparison on classification tasks measured by accuracy.
-Figure 10: Comparison between CP and Tucker decomposition
-V-D2 Experimental results
-First, Figure 9 illustrates that, for most datasets, the two decomposition methods exhibit comparable performance.
-Exceptions include airfoil Self-Noise and electricity datasets, where the two methods show slight performance differences.
-This observation suggests that variations in TD methods do not substantially impact the final results.
-Second, as shown in Figure 10, the CP-based method achieves a 42% reduction in computational time.
+All experiments are repeated 10 times and reported by mean.
+We also report #parameters and #FLOPs to show the computation cost of different methods.
+[…]
+Table 5: Comprehensive Comparison on Benchmark Datasets.
+We evaluate methods on both regression (MSE $\downarrow$) and classification (Accuracy $\uparrow$) tasks.
+To assess efficiency, we also report FLOPs and Params.
+Note that NS-TD achieves comparable or better performance with significantly lower computational costs.
+Dataset MLP KAN NS-SR NS-TD (Ours)
+Perf. FLOPs Param Perf. FLOPs Param Perf. FLOPs Param Perf. FLOPs Param
+Regression Tasks (Metric: MSE $\downarrow$)
+California Housing 0.0890 1568 833 0.1017 19544 7056 0.1135 3104 1601 0.1005 1568 1857
+House Sales 0.0077 752 401 0.0073 10551 3384 0.0075 1488 769 0.0075 2224 1137
+Airfoil 0.0326 152 89 0.0220 3017 684 0.0712 152 89 0.0209 152 393
+Diamonds 0.0053 560 305 0.0040 8241 2520 0.0043 1104 577 0.0045 560 897
+Abalone 0.0249 200 113 0.0232 3740 900 0.0243 392 209 0.0236 584 305
+Bike Sharing 0.0250 656 353 0.0086 6460 1872 0.0081 1296 673 0.0065 1936 993
+Space GA 0.0064 168 97 0.0056 3258 756 0.0063 328 177 0.0056 488 257
+Airlines Delay 0.1605 848 465 0.1563 11997 3816 0.1602 1680 881 0.1560 2512 1297
+Classification Tasks (Metric: Accuracy $\uparrow$)
+Credit 0.7341 608 330 0.7431 3474 864 0.7441 1184 618 0.7471 1760 906
+Heloc 0.7107 432 230 0.7027 7186 1944 0.7117 848 438 0.7077 432 806
+Electricity 0.7762 544 298 0.7962 2992 720 0.7932 1056 554 0.7862 1568 810
+Phoneme 0.7742 448 250 0.8292 6845 2016 0.7942 864 458 0.8242 448 778
+MagicTelescope 0.8444 608 330 0.8479 8770 2736 0.8389 1184 618 0.8495 1184 1226
+Vehicle 0.8106 896 476 0.7576 12138 4032 0.7676 1728 892 0.8176 2560 1308
+Oranges-vs-Grape 0.9105 448 250 0.9302 6845 2016 0.9235 864 458 0.9344 448 250
+Eye Movements 0.5699 400 214 0.5819 6704 1800 0.5899 784 406 0.5940 400 758
+5.4 Comparison of Tensor Decomposition Methods
+Our framework employs the CP decomposition for neuronal discovery. To empirically validate this choice, we conduct a comparative analysis of some tensor decomposition methods used in NS-TD.
+5.4.1 Experimental Settings
+Generally, we follow the previous experimental settings in Section 5.3: 16 tabular datasets and the same network structures as described in Table 5.
+We adopt CP decomposition and Tucker decomposition, respectively, for tensor processing in NS-TD.
+We employ MSE for regression tasks and accuracy for classification tasks.
+5.4.2 Experimental Results
+First, Fig.
+10 illustrates that, for most datasets, the two decomposition methods exhibit comparable performance.
+Exceptions include airfoil Self-Noise and electricity datasets, where two methods show slight performance differences.
+This observation suggests that variations in decompositions for TD methods do not substantially impact the final results.
+Second, as shown in Fig.
+11, the CP-based method achieves a 42% reduction in computational time.
 This result verifies that the CP method offers a computational advantage over the Tucker method.
-VI Competitive Experiments of NeuronSeek-TD over Other Baselines
-VI-A Experiment on Tabular Data
-To further illustrate the superiority of NeuronSeek-TD, we compare it with other state-of-the-art models over 4 real-world tabular tasks: Electron Collision Prediction [^1] https://www.kaggle.com/datasets/fedesoriano/cern-electron-collision-data, Asteroid Prediction [^2] https://www.kaggle.com/datasets/basu369victor/prediction-of-asteroid-diameter, Heart Failure Detection [^3] https://www.kaggle.com/datasets/fedesoriano/heart-failure-prediction/data, and Stellar Classification [^4] https://www.kaggle.com/datasets/fedesoriano/stellar-classification-dataset-sdss17).
-We adopt 7 advanced machine learning methods, namely XGBoost [59], LightGBM [60], CatBoost [61], TabNet [62], TabTransformer [63], FT-Transformer [64], DANETs [65] and NeuronSeek-SR [12] as the baselines to compare with NeuronSeek-TD.
-Moreover, our proposed model and NeuronSeek-SR replace standard neurons with what are discovered by NeuronSeek-TD and NeuronSeek-SR, respectively.
+(a) Regression Task
+(b) Classification Task
+Figure 10: Comparison between CP and Tucker decomposition.
+Figure 11: Time comparison between CP and Tucker decomposition.
+6 Competitive Experiments of NS-TD over Other Baselines
+6.1 Experiment on Tabular Data
+To further illustrate the superiority of NS-TD, we compare it with other state-of-the-art models over 4 real-world tabular tasks: Electron Collision Prediction, Asteroid Prediction, Heart Failure Detection, and Stellar Classification).
+We adopt 7 advanced machine learning methods, namely XGBoost [6], LightGBM [30], CatBoost [21], TabNet [1], TabTransformer [26], FT-Transformer [19], DANETs [5] and NS-SR [17] as the baselines to compare with NS-TD.
+Moreover, our proposed model and NS-SR replace standard neurons with what are discovered by NS-TD and NS-SR, respectively.
 All datasets are randomly split into training, validation, and test sets in a 0.8:0.1:0.1 ratio.
 For regression tasks (electron collision and asteroid prediction), we use the MSE as an evaluation metric, while for classification tasks with imbalanced labels (heart failure detection and stellar classification), we adopt the F1 score.
-Detailed experiment information is shown in Table IV.
-TABLE IV: Details of real-world tabular datasets.
+Detailed experiment information is shown in Table 6.
+Table 6: Details of real-world tabular datasets.
 Dataset Task Data Size Features
@@
 Stellar Classification Classification 100,000 17
-TABLE V: The test results of different models on real-world datasets.
+In Table 7, it can be seen that TabTransformer demonstrates competitive performance on the electron collision dataset.
+However, its effectiveness sharply declines on the asteroid prediction.
+In contrast, CatBoost and TabNet exhibit more consistent yet suboptimal results across four datasets.
+Lastly, the proposed NS-TD is the best performer, achieving the lowest error rates and highest F1 scores uniformly across all four benchmarks.
+Notably, NS-TD significantly outperforms FT-Transformer and DANET.
+This significant margin underscores the efficacy of its task-adaptive architecture in recognizing feature interactions without domain-specific tuning.
+Table 7: The test results of different models on real-world datasets.
 Method Electron Collision (MSE) Asteroid Prediction (MSE) Heart Failure (F1) Stellar Classification (F1)
@@
 DANETs 0.0076 $\pm$ 0.0009 0.1709 $\pm$ 0.1859 0.8948 $\pm$ 0.03 0.9681 $\pm$ 0.002
-NeuronSeek-SR 0.0016 $\pm$ 0.0005 0.0513 $\pm$ 0.0551 0.8874 $\pm$ 0.05 0.9613 $\pm$ 0.002
-NeuronSeek-TD (ours) 0.0011 $\pm$ 0.0003 0.0502 $\pm$ 0.0800 0.9023 $\pm$ 0.03 0.9714 $\pm$ 0.001
-Results are compared in Table V.
-It can be seen that TabTransformer demonstrates competitive performance on the particle collision dataset.
-However, its effectiveness sharply declines on the asteroid prediction.
-In contrast, CatBoost and TabNet exhibit more consistent yet suboptimal results across four datasets.
-Lastly, the proposed NeuronSeek-TD method is the best performer, achieving the lowest error rates and highest F1 scores uniformly across all four benchmarks.
-Notably, NeuronSeek-TD significantly outperforms FT-Transformer and DANET.
-This significant margin underscores the efficacy of its task-adaptive architecture in recognizing feature interactions without domain-specific tuning.
-VI-B Experiments on Images
-To evaluate the performance enhancement of NeuronSeek-TD in computer vision tasks, we integrate it into several established convolutional neural network architectures: ResNet [3], DenseNet [66], SeResNet [67], and GoogLeNet [68].
+NS-SR 0.0016 $\pm$ 0.0005 0.0513 $\pm$ 0.0551 0.8874 $\pm$ 0.05 0.9613 $\pm$ 0.002
+NS-TD (ours) 0.0011 $\pm$ 0.0003 0.0502 $\pm$ 0.0800 0.9023 $\pm$ 0.03 0.9714 $\pm$ 0.001
+6.2 Experiments on Images
+To evaluate the performance enhancement of NS-TD in computer vision tasks, we integrate it into several established convolutional neural network architectures: ResNet [22], DenseNet [24], SeResNet [23], and GoogLeNet [56].
 Notably, we replace the first convolutional layer of each network block with a TN convolutional layer employing the discovered formula $x+\sin x$.
-It should be noted that the original NeuronSeek-SR was not designed for high-dimensional image tasks due to convergence difficulties.
-Our implementation focuses on NeuronSeek-TD and follows the standardized training protocol from a popular image classification repository [^5] https://github.com/weiaicunzai/pytorch-cifar100, incorporating several advanced training techniques: Learning rate scheduling via WarmUpLR and MultistepLR, regularization through Label Smoothing, and optimization using SGD.
+It should be noted that the original NS-SR was not designed for high-dimensional image tasks due to convergence difficulties.
+Our implementation focuses on NS-TD and follows the standardized training protocol from a popular image classification repository [^3] https://github.com/weiaicunzai/pytorch-cifar100, incorporating several advanced training techniques: Learning rate scheduling via WarmUpLR and MultistepLR, regularization through Label Smoothing, and optimization using SGD.
 We use grid search to fine-tune the learning rate from 0.5 to 0.05.
@@
 In particular, images from the STL10 dataset are resized to $32\times 32$ to fit the input of the model.
-The classification results are summarized in Table VI.
-Our experiments demonstrate that task-driven neuronal layers consistently enhance the baseline performance across multiple datasets.
-Notably, NeuronSeek-augmented ResNet18 achieves a 2% accuracy improvement on both CIFAR 100 and STL10 datasets.
-While SeResNet exhibits median performance on CIFAR and SVHN datasets, its performance degrades on STL10; however, TN layers compensate for this degradation by boosting its accuracy by 10%.
-These results suggest that task-driven neuronal formula discovery can effectively unlock the latent potential of fundamental network architectures.
-TABLE VI: Accuracy of compared methods on small-scale image datasets.
+The classification results are summarized in Table 8.
+Table 8: Accuracy of compared methods on small-scale image datasets.
 CIFAR10 CIFAR100 SVHN STL10
@@
 NeuronSeek-GoogleNet 0.9400 0.7519 0.9658 0.7298
-VII Discussions and Conclusions
+Our experiments demonstrate that task-driven neuronal layers consistently enhance the baseline performance across multiple datasets.
+Notably, NeuronSeek-augmented ResNet18 achieves a 2% accuracy improvement on both CIFAR 100 and STL10 datasets.
+While SeResNet exhibits median performance on CIFAR and SVHN datasets, its performance degrades on STL10; however, TN layers compensate for this degradation by boosting its accuracy by 10%.
+These results suggest that task-driven neuronal formula discovery can effectively unlock the latent potential of fundamental network architectures.
+7 Conclusions
 In this paper, we have proposed a novel method in the framework of NeuronSeek to enhance the capabilities of a neural network.
-Compared to neural architecture search (NAS) [72], NeuronSeek provides a different approach.
-Rather than modifying the overall model structure, NeuronSeek optimizes solely at the neuronal level by searching for optimal aggregation formulas.
+Compared to neural architecture search (NAS) [61], NeuronSeek emphasizes the importance of neurons by searching for optimal aggregation formulas.
 Firstly, we have introduced tensor decomposition to prototype stable task-driven neurons.
 Secondly, we have provided theoretical guarantees demonstrating that these neurons can maintain the super-super-expressive property.
-At last, through extensive empirical evaluation, we show that our proposed framework consistently outperforms baseline models while achieving the state-of-the-art results across diverse datasets and tasks.
-These advancements establish task-driven neurons as an effective addition in the design of task-specific neural network components, offering both theoretical soundness and practical efficacy.
+At last, through extensive empirical evaluation, we have shown that our proposed framework consistently outperforms baseline models while achieving the state-of-the-art results across diverse datasets and tasks.
+These advancements establish task-driven neurons as an effective addition in the design of task-specific neural network components, offering both theoretical soundness and practical efficacy and stability.
+Appendix A Supplementary Results
+A.1 Experiments on Synthetic Data
+Supplementary results on synthetic data (Section 5.1) are illustrated as follows.
+Table 9: The reconstruction accuracy ($\text{Mean}_{\pm\text{SEM}}$) of pure mode.
+Method Pure Mode ($\sum x^{k}$)
+10 30 50 100
+SR (PySR) $1612_{0.0224} 1.1$$1493_{0.0061}$ $1.2325_{0.0042}$ $1.2481_{0.0093}$
+EQL $0.4524_{0.0232}$ $0.5342_{0.0141}$ $0.5705_{0.0123}$ $0.6923_{0.0132}$
+NS-SR (VSR) $0.2324_{0.0041}$ $0.2541_{0.0092}$ $0.2825_{0.0063}$ $0.2953_{0.0071}$
+MetaSymNet $0.8324_{0.0642}$ $0.8541_{0.0053}$ $0.9825_{0.0081}$ $0.9843_{0.0092}$
+NS-TD $\mathbf{0.0543}_{0.0062}$ $\mathbf{0.0632}_{0.0051}$ $\mathbf{0.0854}_{0.0083}$ $\mathbf{0.0891}_{0.0082}$
+Table 10: The reconstruction accuracy ($\text{Mean}_{\pm\text{SEM}}$) of interact mode.
+Method Interact Mode (Vector Interact)
+10 30 50 100
+SR (PySR) $0.8754_{0.0102}$ $1.0042_{0.0084}$ $0.9776_{0.0192}$ $0.9763_{0.0091}$
+EQL $0.3875_{0.0174}$ $0.4862_{0.0021}$ $0.5614_{0.0042}$ $0.4521_{0.0073}$
+NS-SR (VSR) $0.6972_{0.0053}$ $0.7734_{0.0052}$ $0.7326_{0.0074}$ $0.8145_{0.0042}$
+MetaSymNet $0.8632_{0.0064}$ $1.0804_{0.0093}$ $1.0193_{0.0102}$ $0.9942_{0.0073}$
+NS-TD $\mathbf{0.0512}_{0.0041}$ $\mathbf{0.0583}_{0.0082}$ $\mathbf{0.0724}_{0.0053}$ $\mathbf{0.0945}_{0.0042}$
+Table 11: The reconstruction accuracy ($\text{Mean}_{\pm\text{SEM}}$) of hybrid mode.
+Method Hybrid Mode (Mixture)
+10 30 50 100
+SR (PySR) $1.0635_{0.0083}$ $0.9821_{0.0072}$ $0764_{0.0123} 1.1$$1018_{0.0241}$
+EQL $0.3126_{0.0052}$ $0.4113_{0.0051}$ $0.4237_{0.0063}$ $0.4472_{0.0061}$
+NS-SR (VSR) $0.7783_{0.0081}$ $0.7824_{0.0072}$ $0.8281_{0.0031}$ $0.8685_{0.0083}$
+MetaSymNet $1.1345_{0.0124}$ $0.9526_{0.0082}$ $1093_{0.0173} 1.1$$1334_{0.0162}$
+NS-TD $\mathbf{0.0423}_{0.0041}$ $\mathbf{0.0621}_{0.0043}$ $\mathbf{0.0735}_{0.0071}$ $\mathbf{0.1112}_{0.0012}$
+A.2 Experiments on Tabular Data
+We present the supplementary results of Section 5.3 to showcase the superiority of the proposed NS-TD.
+Table 12: The supplementary results with standard deviations on the tabular datasets. The metrics are presented in the format $\text{Mean}_{\text{Std}}$.
+Dataset MLP KAN NS-SR NS-TD (Ours)
+Regression Tasks (Metric: MSE $\downarrow$)
+California Housing $\mathbf{0.0890}_{0.0021}$ $0.1017_{0.0035}$ $0.1135_{0.0052}$ $\underline{0.1005}_{0.0050}$
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Airlines Delay $0.1605_{0.0050}$ $\underline{0.1563}_{0.0055}$ $0.1602_{0.0055}$ $\mathbf{0.1560}_{0.0052}$
+Classification Tasks (Metric: Accuracy $\uparrow$)
+Credit $0.7341_{0.0085}$ $0.7431_{0.0102}$ $\underline{0.7441}_{0.0092}$ $\mathbf{0.7471}_{0.0092}$
+Heloc $\underline{0.7107}_{0.0120}$ $0.7027_{0.0135}$ $\mathbf{0.7117}_{0.0145}$ $0.7077_{0.0113}$
+Electricity $0.7762_{0.0070}$ $\mathbf{0.7962}_{0.0085}$ $\underline{0.7932}_{0.0075}$ $0.7862_{0.0087}$
+Phoneme $0.7742_{0.0210}$ $\mathbf{0.8292}_{0.0245}$ $0.7942_{0.0207}$ $\underline{0.8242}_{0.0252}$
+MagicTelescope $0.8444_{0.0088}$ $\underline{0.8479}_{0.0095}$ $0.8389_{0.0092}$ $\mathbf{0.8495}_{0.0079}$
+Vehicle $\underline{0.8106}_{0.0340}$ $0.7576_{0.0380}$ $0.7676_{0.0362}$ $\mathbf{0.8176}_{0.0354}$
+Oranges-vs-Grape $0.9105_{0.0150}$ $\underline{0.9302}_{0.0210}$ $0.9235_{0.0160}$ $\mathbf{0.9344}_{0.0240}$
+Eye Movements $0.5699_{0.0135}$ $0.5819_{0.0155}$ $\underline{0.5899}_{0.0145}$ $\mathbf{0.5940}_{0.0164}$